### v1

# Strong convergence of uniformly random permutation representations of surface groups

Abstract
Let $\Gamma$ be the fundamental group of a closed orientable surface of genus at least two. Consider the composition of a uniformly random element of $\mathrm{Hom}(\Gamma,S_{n})$ with the $(n-1)$-dimensional irreducible representation of $S_{n}$. We prove the strong convergence in probability as $n\to\infty$ of this sequence of random representations to the regular representation of $\Gamma$.
As a consequence, for any closed hyperbolic surface $X$, with probability tending to one as $n\to\infty$, a uniformly random degree-$n$ covering space of $X$ has near optimal relative spectral gap — ignoring the eigenvalues that arise from the base surface $X$.
To do so, we show that the polynomial method of proving strong convergence can be extended beyond rational settings.
To meet the requirements of this extension we prove two new kinds of results. First, we show there are effective polynomial approximations of expected values of traces of elements of $\Gamma$ under random homomorphisms to $S_{n}$. Secondly, we estimate the growth rates of probabilities that a finitely supported random walk on $\Gamma$ is a proper power after a given number of steps.
1 Introduction
In the following $\Gamma=\Gamma_{g}$ will be the fundamental group of a closed orientable surface of genus $g\geq 2$. We view $g$ and $\Gamma$ as fixed in the paper. We denote by $S_{n}$ the group of permutations of $[n]\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\{1,\ldots,n\}$ and $\mathcal{U}(n)$ the group of $n\times n$ complex unitary matrices. The $n$-dimensional linear representation of $S_{n}$ by permutation matrices is not irreducible but has an $(n-1)$-dimensional irreducible subrepresentation
$\mathrm{std}:S_{n}\to\mathcal{U}(n-1)$
obtained by removing the non-zero invariant vectors. The group $\Gamma$ also has an important unitary representation called the regular representation:
$\lambda:\Gamma\to\mathcal{U}(\ell^{2}(\Gamma)),\quad\lambda(\gamma)[f](x)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}f(\gamma^{-1}x).$
For $n\in\mathbf{N}$ the set of homomorphisms
$\mathbb{X}_{g,n}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathrm{Hom}(\Gamma,S_{n})$
is finite and we equip it with the uniform probability measure. Let $\mathbf{C}[\Gamma]$ denote the group algebra of $\Gamma$. Every linear representation of $\Gamma$ extends linearly to one of $\mathbf{C}[\Gamma]$.
Theorem 1.1.
For all $x\in\mathbf{C}[\Gamma]$, for $\phi_{n}\in\mathrm{Hom}(\Gamma,S_{n})$ uniformly random, as $n\to\infty$
$\|\mathrm{std}\circ\phi_{n}(x)\|\to\|\lambda(x)\|$
in probability. The norms on each side are operator norms.
In the language of strong convergence of unitary representations, Theorem 1.1 says precisely that the random representations $\mathrm{std}\circ\phi_{n}$ converge strongly in probability to $\lambda$. See [Mag25] for a survey on strong convergence. Theorem 1.1 is the analog, for surface groups, of the main theorem of Bordenave–Collins [BC19] for free groups. In §1.4 we cover the history of strong convergence results prior to Theorem 1.1.
Random hyperbolic surfaces. One of the primary motivations for Theorem 1.1 is a known important consequence on the spectral theory of random hyperbolic surfaces. A hyperbolic surface $X$ is a complete Riemannian surface of constant curvature $-1$. Such a surface has an associated Laplacian operator $\Delta_{X}$. If $X$ is closed, the spectrum of $\Delta_{X}$ as an unbounded operator on $L^{2}(X)$ is discrete consisting of real eigenvalues that accumulate only at infinity. If $X$ is connected, the bottom of the spectrum of $\Delta_{X}$ is a simple eigenvalue at 00 of multiplicity one.
The first non-zero eigenvalue, $\lambda_{1}(X)$, is of fundamental significance to e.g. the dynamics of the geodesic flow on $X$, the statistics of lengths of closed geodesics, and notions of connectedness of $X$. In cases where $X$ is a congruence hyperbolic surface, Selberg recognized [Sel65] that the condition $\lambda_{1}(X)\geq\frac{1}{4}$ — the Selberg Eigenvalue conjecture — is the archimedean analog of the Ramanujan–Petersson conjectures from number theory, and from the point of view of the Selberg zeta function is also the analog of the Riemann hypothesis for these surfaces.
If $\{X_{n}\}_{n=1}^{\infty}$ are a sequence of closed hyperbolic surfaces with genera tending to $\infty$, then $\limsup_{n\to\infty}\lambda_{1}(X_{n})\leq\frac{1}{4}$ [Hub74] and hence $\frac{1}{4}$ is the asymptotically optimal value of $\lambda_{1}(X_{n})$ for such surfaces. The number $\frac{1}{4}$ arises here as the bottom of the spectrum of the Laplacian on hyperbolic space of two dimensions.
Theorem 1.2.
Let $X$ be a closed connected hyperbolic surface. Let $X_{n}$ denote a uniformly random degree $n$ covering space of $X$ [^1] If $\pi_{1}(X)\cong\Gamma$, then by definition, a uniformly random degree-$n$ cover of $X$ is given by a uniformly random element of $\mathbb{X}_{g,n}$ via the bijection between $\mathbb{X}_{g,n}$ and covers with a labeled fiber; see [MP23, §1]. With probability tending to one as $n\to\infty$
$\mathrm{spec}(\Delta_{X_{n}})\cap\left[0,\frac{1}{4}-o(1)\right)=\mathrm{spec}(\Delta_{X})\cap\left[0,\frac{1}{4}-o(1)\right),$
with the same multiplicities on both sides.
In particular, if $\lambda_{1}(X)\geq\frac{1}{4}$, then $\lambda_{1}(X_{n})\to\frac{1}{4}$ in probability as $n\to\infty$.
We omit the proof that Theorem 1.2 follows from Theorem 1.1, which is the same as in [LM25, Appendix A]. [^2] This passage as presented is non-effective. For an effective version one can see [Hid25].
The existence of covering spaces of hyperbolic surfaces that have optimal spectral gaps was proved in [HM23, LM25]. What Theorem 1.2 shows is that not only do such surfaces exist, but that this is in fact the typical behavior of covering spaces of closed hyperbolic surfaces. In §1.4 we describe prior results to Theorem 1.2.
1.1 Technical innovation I: The generalized polynomial method
The polynomial method, which was introduced in the recent work of Chen, Garza-Vargas, Tropp, and the third named author [CGVTvH24, CGVvH24] and further developed by de la Salle and the first named author [MdlS24b], provides a powerful approach for establishing strong convergence in situations that are outside the reach of previous methods. To date, however, all applications of the polynomial method have relied fundamentally on the property that the spectral statistics of the random matrices that arise in these applications are rational functions of $\frac{1}{n}$. This restrictive property fails manifestly for the model considered in this paper. [^3] Indeed, in the setting of this paper, the spectral statistics are not even analytic as a function of $\frac{1}{n}$. This can be verified by checking, by an explicit computation, that the coefficients $a_{i}(\gamma)$ in Theorem 1.6 grow faster than exponentially in $i$ when $\gamma$ is one of the standard generators of $\Gamma$. We omit the details.
The basis for the present paper is a refinement of the polynomial method, which makes it applicable to much more general situations where the dimension dependence of the spectral statistics need not even be analytic. This greatly expands the potential range of applications of this method. More precisely, this method will give rise to the following general criterion for strong convergence.
Let $\Lambda$ be a finitely generated group with a finite generating set $S$. For $\gamma\in\Lambda$, denote by $|\gamma|$ the word length of $\gamma$ with respect to $S$. Denote by $\lambda$ the left regular representation of $\Lambda$, and for each $n\in\mathbf{N}$ let
$\pi_{n}:\Lambda\to\mathcal{U}(n)$
be a random unitary representation. We denote by $\mathop{\mathrm{Tr}}M$ the trace of a matrix $M$ and by $\mathop{\mathrm{tr}}M$ its normalized trace. We let $\tau$ be the canonical trace on the reduced group $C^{*}$-algebra $C_{{\rm red}}^{*}(\Lambda)$. (The definitions of these notions are recalled in §1.6.)
To establish that $\pi_{n}\to\lambda$ strongly in probability, we will need two assumptions. The first is an effective $\frac{1}{n}$-expansion for the expected spectral statistics.
Assumption 1.3.
For every $\gamma\in\Lambda$, there exist $u_{k}(\gamma)$ with $u_{0}(\gamma)=\tau(\lambda(\gamma))$ and
$\Bigg{|}\mathbb{E}\big{[}{\mathop{\mathrm{tr}}\pi_{n}(\gamma)}\big{]}-\sum_{k=0}^{q-1}\frac{u_{k}(\gamma)}{n^{k}}\Bigg{|}\leq\frac{(Cq)^{Cq}}{n^{q}}$
for all $q\geq|\gamma|$ and $n\geq Cq^{C}$, where $C\geq 1$ is a constant.
In the following, we will extend $u_{k}$ linearly to the group algebra $\mathbf{C}[G]$. The second assumption is that $u_{1}$ is “tempered” in the sense of [CGVTvH24, MdlS24b].
Assumption 1.4.
For every self-adjoint $x\in\mathbf{C}[\Lambda]$, we have
$\limsup_{p\to\infty}|u_{1}(x^{p})|^{1/p}\leq\|\lambda(x)\|.$
The following theorem is proved in §3.
Theorem 1.5.
Suppose that Assumptions 1.3 and 1.4 hold. Then
$\|\pi_{n}(x)\|\leq\|\lambda(x)\|+o(1)\quad\text{with probability}\quad 1-o(1)$
as $n\to\infty$ for every $x\in\mathbf{C}[\Lambda]$.
Theorem 1.5 gives a strong convergence upper bound. If $C_{{\rm red}}^{*}(\Lambda)$ has a unique trace, this automatically implies the lower bound [MdlS24b, §5.3]. We state the result here in a qualitative form for simplicity, but Theorem 3.3 below in fact provides a quantitative form of this result with a polynomial convergence rate.
Let us highlight two important features of the above result.
  • Previous applications of the polynomial method required that $\mathbb{E}[\mathop{\mathrm{tr}}\pi_{n}(\gamma)]$ is a rational function of $\frac{1}{n}$. Assumption 1.3 is a drastic relaxation of this assumption: it requires only that the spectral statistics are in a Gevrey class (so they need not even be analytic, let alone rational) as a function of $\frac{1}{n}$.
  • To verify Assumption 1.3, it is only necessary to consider individual group elements $\gamma\in\Lambda$ rather than elements of the group algebra $x=\sum_{\gamma}\alpha_{\gamma}\gamma\in\mathbf{C}[\Lambda]$.
The latter is crucial for applications. In principle, one may expect that a method that applies to polynomial (or rational) spectral statistics could be adapted to spectral statistics that are well approximated by polynomials. This is what Assumption 1.3 ensures. However, a key feature of the polynomial method is that it captures cancellations between the terms of $x\in\mathbf{C}[\Lambda]$ in the $\frac{1}{n}$-expansion of $\mathbb{E}[\mathop{\mathrm{tr}}\pi_{n}(x)]$ [CGVTvH24, §2.3]. It is not clear that such cancellations are preserved when the spectral statistics are approximated by polynomials. Since we do not know how to establish a priori bounds that capture cancellations, it is essential that this is not required by Assumption 1.3. We elaborate further on this point in the remarks at the end of §3.
1.2 Technical innovation II: Effective polynomial approximation
We aim to apply the criterion for strong convergence from Theorem 1.5 to the setting that $\Lambda=\Gamma=\Gamma_{g}$ is a surface group with $g\geq 2$, and
$\pi_{n}=\mathrm{std}\circ\phi_{n},$
where $\phi_{n}$ is a uniformly distributed random element of $\mathrm{Hom}(\Gamma_{g},S_{n})$. Thus we need to establish Assumptions 1.3 and 1.4 for this setting. We begin with Assumption 1.3.
For $\sigma\in S_{n}$ let $\mathsf{fix}(\sigma)$ denote the number of fixed points of $\sigma$. For a random variable $f$ on $\mathbb{X}_{g,n}$ we denote by $\mathbb{E}_{g,n}[f]$ the expectation of $f$ with respect to the uniform measure. Natural random variables on $\mathbb{X}_{g,n}$ come from elements of $\gamma$: given $\gamma\in\Gamma$ we define
$\mathsf{fix}_{\gamma}$ $:\mathbb{X}_{g,n}\to\mathbf{Z}$
$\mathsf{fix}_{\gamma}$ $:\phi\mapsto\mathsf{fix}(\phi(\gamma)).$
Then we have
$\mathbb{E}[\mathop{\mathrm{Tr}}\pi_{n}(\gamma)]=\mathbb{E}_{g,n}[\mathsf{fix}_{\gamma}]-1.$ (1.1)
An asymptotic expansion for $\mathbb{E}_{g,n}[\mathsf{fix}_{\gamma}]$ was obtained in the previous work [MP22, MP23] of the first two authors.
Theorem 1.6 ([MP23, Thm. 1.1]).
For each $\gamma\in\Gamma$ there exists a (unique) sequence $\left\{a_{i}(\gamma)\right\}_{i=-1}^{\infty}$ such that for any $q\in\mathbf{N}$, as $n\to\infty$
$\mathbb{E}_{g,n}[\mathsf{fix}_{\gamma}]=a_{-1}(\gamma)n+a_{0}(\gamma)+a_{1}(\gamma)n^{-1}+\cdots+a_{q-1}(\gamma)n^{-(q-1)}+O_{q,\gamma}\left(n^{-q}\right),$
where the implied constant in the big O depends on $q$ and $\gamma$.
In §4, we prove the following theorem that strengthens Theorem 1.6. Here and in the sequel, we fix a standard set of generators of $\Gamma$ of size $2g$ coming from the presentation
$\Gamma=\left\langle a_{1},b_{1},\ldots,a_{g},b_{g}\,\mid\,[a_{1},b_{1}]\cdots[a_{g},b_{g}]\right\rangle,$ (1.2)
and for $\gamma\in\Gamma$ write $|\gamma|$ for the word length of $\gamma$ with respect to these generators.
Theorem 1.7.
There exists $C=C(g)>1$ such that for all $q,n\in\mathbf{N}$ with $n\geq Cq^{C}$, and for all $\gamma\in\Gamma$ with $|\gamma|\leq q$,
$\left|\mathbb{E}_{g,n}[\mathsf{fix}_{\gamma}]-a_{-1}(\gamma)n-a_{0}(\gamma)-a_{1}(\gamma)n^{-1}-\cdots-a_{q-1}(\gamma)n^{-(q-1)}\right|\leq(Cq)^{Cq}n^{-q}.$
The proof of this effective asymptotic expansion proceeds in two parts. The first part of the analysis relies on an exact expansion of $\mathbb{E}_{g,n}[\mathsf{fix}_{\gamma}]$ in terms of all irreducible representations of $S_{n}$ that is developed in [MP23]. A careful analysis shows that we may truncate this expansion only to those representations defined by Young diagrams $\lambda\vdash n$ with $O(q)$ boxes outside the first row and column, with an effective error bound as in Assumption 1.3. This is Proposition 4.4.
It is not clear, however, how to analyze the truncated expansion using the methods of [MP23]. Instead, we introduce a new method for analyzing the truncated expansion by means of a combinatorial integration formula for the symmetric group. This enables us to show that the truncated expansion is in fact a rational function of $\frac{1}{n}$. An effective error bound for the truncated expansion can be then achieved using the analytic theory of polynomials, in a manner closely analogous to the polynomial method for random matrix models with rational spectral statistics [CGVTvH24, CGVvH24, MdlS24b].
1.3 Technical innovation III: Geometry of proper powers
It remains to establish Assumption 1.4 for surface groups.
For every element $1\neq\gamma\in\Gamma$, there is a unique $\gamma_{0}\in\Gamma$ and $k\in\mathbf{N}$ with $\gamma=\gamma_{0}^{\,k}$ such that $\gamma_{0}$ is not a proper power. [^4] This follows immediately from the fact the non-trivial elements of $\Gamma$ admit cyclic centralizer. The latter fact can be found, e.g., in [FM11, p. 23]. We denote by $\Gamma_{\rm np}$ the set of non-powers in $\Gamma$. Denote by $\omega(k)$ the number of positive divisors of $k$, and by abuse of notation, we define
$\omega(\gamma)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\begin{cases}0&\text{if }\gamma=1,\\ \omega(k)&\text{if }\gamma=\gamma_{0}^{k}\text{ with }\gamma_{0}\in\Gamma_{\rm np}.\end{cases}$
In the notation of Theorem 1.7, it is shown in [MP23, Thm. 1.2] that $a_{0}(\gamma)=\omega(\gamma)$. Thus $u_{1}$ in Assumption 1.4 can be expressed for any $x=\sum_{\gamma\in\Gamma}\alpha_{\gamma}\gamma\in\mathbf{C}[\Gamma]$ as
$u_{1}(x)=\sum_{\gamma}\alpha_{\gamma}\left[\omega(\gamma)-1\right]$ (1.3)
using (1.1), see §6. Note that $u_{1}$ is determined only by the contributions of $\gamma\in\Gamma$ that are either the identity or a proper power, since $\omega(\gamma)-1=0$ for $\gamma\in\Gamma_{\rm np}$.
We prove the following theorem in §5.
Theorem 1.8.
For any self adjoint $x\in\mathbf{C}[\Gamma]$,
$\limsup_{p\to\infty}\left|u_{1}(x^{p})\right|^{1/p}\leq\left\|\lambda(x)\right\|.$
To prove Theorem 1.8 it suffices, by [MdlS24b, Prop. 6.3] and as surface groups satisfy the rapid decay property [Jol90, Thm. 3.2.1], to control the exponential growth rate of the probability that a random walk in $\Gamma$ with an arbitrary finite set of generators lands on a proper power.
In the free group setting, the latter is readily accomplished by a simple spectral argument that dates back to [Fri03, Lem. 2.4]. This argument relies on the following elementary property of free groups: if a word $w=a_{i_{1}}\cdots a_{i_{q}}$ in the free generators $a_{1},\ldots,a_{r}$ reduces to a proper power $\gamma_{0}^{k}$ and we write $\gamma_{0}=bhb^{-1}$ where $h$ is the cyclic reduction of $\gamma_{0}$, then $w$ must be a concatenation of five consecutive sub-words that reduce to $b,h,h,h^{k-2}$, and $b^{-1}$, respectively.
The difficulty in the setting of surface groups is that such a property is no longer true. However, we will show in Lemma 5.6 that an approximate form of this property remains valid: if a word $w$ in the standard generators of $\Gamma$ is equivalent in $\Gamma$ to a proper power $\gamma_{0}^{k}$, then there must exist $b,h\in\Gamma$ so that $w$ is a concatenation of five consecutive sub-words which are equivalent in $\Gamma$ to elements that are $c\log|w|$-close to $b,h,h,h^{k-2}$, and $b^{-1}$ (here $u$ is $\ell$-close to $v$ if $u$ is equivalent in $\Gamma$ to $\gamma v\gamma^{\prime}$ with $|\gamma|,|\gamma^{\prime}|\leq\ell$). The proof of this fact makes crucial use of the hyperbolic geometry of $\mathrm{Cay}(\Gamma)$. This property suffices to conclude the proof of Assumption 1.4 for surface groups.
1.4 Prior results
Strong convergence
The first result in the spirit of Theorem 1.1 was the breakthrough result of Haagerup and Thorbjørnsen [HT05], who proved that i.i.d. GUE random matrices strongly converge to free semicircular random variables. Let $\mathbb{\mathbb{\mathbf{F}}}_{r}$ denote a free group of rank $r$. Collins and Male proved in [CM14] that Haar-distributed elements of $\mathrm{Hom}(\mathbf{F}_{r},\mathcal{U}(n))\cong\mathcal{U}(n)^{r}$ strongly converge to the regular representation of $\mathbb{\mathbb{\mathbf{F}}}_{r}$. In another breakthrough [BC19], Bordenave and Collins proved that uniformly random elements of $\mathrm{Hom}(\mathbb{\mathbb{\mathbf{F}}}_{r},S_{n})$ composed with $\mathrm{std}$ strongly converge to the regular representation of $\mathbb{\mathbb{\mathbf{F}}}_{r}$.
Subsequent developments include results on higher dimensional representations of $\mathcal{U}(n)$ or $S_{n}$ [BC24b, CGVTvH24, MdlS24b, Cas25], results related to Hayes’ work [Hay22] on the Peterson-Thom conjecture [BC22, BC24a, Par24, CGVvH24], and work of Austin on annealed almost periodic entropy [Aus24].
Regarding the existence of finite dimensional unitary representations that strongly converge to the regular representation, results have been obtained for various classes of groups [LM25, MdlS24a, MT23] other than free groups. Most relevant to the current paper is the result of the first named author and Louder [LM25, Cor. 1.2] that for surface groups $\Gamma$ as in this paper, there exists a sequence of $\{\phi_{n}\in\mathrm{Hom}(\Gamma,S_{n})\}_{n=1}^{\infty}$ such that $\mathrm{std}\circ\phi_{n}$ strongly converge to $\lambda$. Theorem 1.1 shows not only that such homomorphisms exist, but that this is in fact the typical behavior.
The recent works [CGVTvH24, CGVvH24] introduced a new approach to strong convergence, called the polynomial method, that is based almost entirely on soft arguments, in contrast to previous works that were based on problem-specific analytic methods or delicate combinatorial estimates. This method has made it possible to establish strong convergence in previously inaccessible situations and to achieve improved quantitative results. These works form the basis for the generalized polynomial method that is developed in the present paper (see §1.1 and §3).
Hyperbolic surfaces
Whether there exists a sequence of closed hyperbolic surfaces $\{X_{n}\}_{n=1}^{\infty}$ with genera tending to $\infty$ such that $\lambda_{1}(X_{n})\to\frac{1}{4}$ was an old question of Buser [Bus84]. In view of analogous results on optimal spectral gaps of random regular graphs [Alo86, Fri08, Pud15, Bor20] (for the state-of-the-art see the recent breakthrough [HMY25]), it is natural to approach this question through the study of random hyperbolic surfaces. We presently discuss previous results on three models of hyperbolic surfaces:
  1. 1.
    The Brooks–Makover (BM) model and variations, obtained by constructing a random cusped hyperbolic surface out of gluing $m$ ideal triangles and then compactifying the cusps.
  2. 2.
    The Weil–Petersson (WP) random model that comes from the Weil–Petersson Kähler form on the moduli space of genus $g$ hyperbolic surfaces.
  3. 3.
    Models of random degree $n$ covering spaces.
We say events hold w.h.p. (with high probability) if they hold with probability tending to one as the relevant parameter $(m/g/n)$ tends to infinity. In random cover models we understand that $\lambda_{1}$ is measured ignoring the eigenvalues of the base surface.
The first result that random closed hyperbolic surfaces in the BM model have a uniform positive lower bound on $\lambda_{1}$ was due to Brooks–Makover [BM04]. The same result for WP random surfaces was obtained by Mirzakhani in [Mir13]. The explicit bound $\lambda_{1}\geq\frac{3}{16}-o(1)$ w.h.p. for closed surfaces was first obtained in [MNP22] for the uniform random covering model. Following this, the same result was obtained in the WP model by Wu–Xue [WX22] and Lipnowski–Wright [LW24]. The latter was improved to $\lambda_{1}\geq\frac{2}{9}-o(1)$ w.h.p. by Anantharaman–Monk [AM23].
The optimal bound $\lambda_{1}\geq\frac{1}{4}-o(1)$ w.h.p. was finally achieved by Hide and the first named author [HM23] in a random cover model for non-closed hyperbolic surfaces. This suffices, by means of a compactification procedure, to settle Buser’s conjecture in the affirmative. The same proof applies verbatim to a model of gluing hyperbolic triangles [^5] Although it was not pointed out at the time, the proof of [HM23] works just as well for random covers of the quotient of $\mathbb{H}^{2}$ by reflections in the sides of an ideal hyperbolic triangle. In this case, the input is the result of Bordenave and Collins for random homomorphisms from the free product $\mathbf{Z}/2\mathbf{Z}*\mathbf{Z}/2\mathbf{Z}*\mathbf{Z}/2\mathbf{Z}$ to $S_{n}$. This is not exactly the same as the Brooks–Makover gluing, however. and as a result, one obtains $\lambda_{1}\geq\frac{1}{4}-o(1)$ w.h.p. in close cousins of the BM model for closed surfaces. In [LM25], the existence of covers of closed hyperbolic surfaces such that $\lambda_{1}\geq\frac{1}{4}-o(1)$ was proved without a compactification procedure.
Very recently, it was shown in the impressive work of Anantharaman–Monk [AM25] that $\lambda_{1}\geq\frac{1}{4}-o(1)$ w.h.p. in the WP model of closed surfaces. This shows that the optimal spectral gap property is in fact the typical behavior, in the sense of the Weil-Petersson measure, of closed hyperbolic surfaces as $g\to\infty$.
Despite the above advances, the question whether optimal spectral gaps $\lambda_{1}\geq\frac{1}{4}-o(1)$ are also typical for random covering spaces has remained open. The reason is that strong convergence of uniform random homomorphisms of surface groups to $S_{n}$ was outside the reach of previous methods. Thus [HM23, LM25] had to rely on reductions to the case of free groups: either by working with non-compact surfaces, or by embedding surface groups into an iterated sequence of extensions of centralizers of a free group which results in a complicated randomization that is far from uniform. The reason we can resolve this question here (Theorem 1.2) is that the new methods introduced in this paper resolve the corresponding strong convergence problem (Theorem 1.1). [^6] We emphasize that the methods of the present paper are completely independent of those used in [MNP22] to obtain the suboptimal bound $\lambda_{1}\geq\frac{3}{16}-o(1)$ for the uniform random cover model, with the exception of the self-contained Proposition 4.1 that appears in [MNP22, Prop. 4.6].
1.5 Paper organization
The remainder of this paper is organized as follows. In §2, we state some basic inequalities for polynomials and rational functions that will be used in the sequel. The proofs of Theorems 1.5, 1.7, and 1.8 are subsequently developed in §3, §4, and §5, respectively. Finally, we assemble the above ingredients in §6 to complete the proof of Theorem 1.1.
1.6 Notation
We write $\mathbf{N}=\{1,2,3,\ldots\}$, $\mathbf{Z}_{+}=\{0,1,2,\ldots\}$, and $\mathbf{Z}=\{\ldots,-1,0,1,\ldots\}$. For $n\in\mathbf{N}$ we write $[n]=\{1,\ldots,n\}$. The falling Pochhammer symbol and double factorial are
$(n)_{m}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}n(n-1)\cdots(n-m+1),\qquad\quad(2n-1)!!\stackrel{{\scriptstyle\text{def}}}{{=}}(2n-1)(2n-3)\cdots 3\cdot 1.$
A set partition of $[n]$ is a collection of disjoint subsets whose union is $[n]$. We write $\mathrm{Part}([n])$ for the set partitions of $[n]$.
For $f,g:\mathbf{N}\to\mathbf{R}$, we write $f=O(g)$ if there is a universal constant $C>0$ — depending on no other parameters — such that for all $n\in\mathbf{N},$ $|f(n)|\leq Cg(n)$. If $C$ depends on other parameters this is indicated with a subscript to the big $O$.
We denote by $f^{(m)}$ the $m$th derivative of a univariate function $f$, and denote by $\|f\|_{[a,b]}=\sup_{z\in[a,b]}|f(z)|$ the uniform norm on the interval $[a,b]$.
Groups and $C^{*}$-algebras
Let $\Lambda$ be a finitely generated group with a finite generating set $S$. We denote by $|\gamma|$ the word length of $\gamma\in\Lambda$ with respect to $S$.
The group algebra $\mathbf{C}[\Lambda]$ has a natural involution $(\sum_{\gamma}\alpha_{\gamma}\gamma)^{*}=\sum_{\gamma}\bar{\alpha}_{\gamma}\gamma^{-1}$. Thus we can speak of self-adjoint elements of the group algebra; these become self-adjoint operators in any unitary representation. For any univariate polynomial $h$ and $x\in\mathbf{C}[\Lambda]$, we interpret $h(x)$ in the obvious algebraic sense. We denote by $|x|$ the maximum word length of a group element in the support of $x$. We define
$\|x\|_{C^{*}(\Lambda)}=\sup_{\pi}\|\pi(x)\|$
for $x\in\mathbf{C}[\Lambda]$, where the supremum is taken over all unitary representations of $\Lambda$.
Let $\lambda$ be the regular representation of $\Lambda$, and extend it linearly to $\mathbf{C}[\Lambda]$. The reduced $C^{*}$-algebra $C^{*}_{\rm red}(\Lambda)$ of $\Lambda$ is the norm-closure of $\{\lambda(x):x\in\mathbf{C}[\Lambda]\}$.
For $M\in\mathrm{M}_{n}(\mathbf{C})$, let $\mathop{\mathrm{Tr}}M$ be the trace and $\mathop{\mathrm{tr}}M=\frac{1}{n}\mathop{\mathrm{Tr}}M$ be the normalized trace. We denote by $\tau$ the canonical trace on $C^{*}_{\rm red}(\Lambda)$, i.e., $\tau(a)=\langle\delta_{e},a\delta_{e}\rangle$, and by
$\|a\|_{L^{2}(\tau)}=\tau(a^{*}a)^{1/2}$
for $a\in C^{*}_{\rm red}(\Lambda)$. In particular, $\|\lambda(x)\|_{L^{2}(\tau)}=(\sum_{\gamma}|\alpha_{\gamma}|^{2})^{1/2}$ for $x=\sum_{\gamma}\alpha_{\gamma}\gamma$.
Acknowledgments
We thank Bram Petri and Joe Thomas for discussions related to this work.
This work was conducted while D.P. and R.v.H. were members of the Institute of Advanced Study in Princeton, NJ, which is gratefully acknowledged for providing a fantastic mathematical environment.
Funding (M.M.): This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 949143).
Funding (D.P.): This work was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 850956), by the Israel Science Foundation, ISF grants 1140/23, as well as National Science Foundation under Grant No. DMS-1926686.
Funding (R.v.H.): This work was supported in part by NSF grant DMS-2347954.
2 Analysis of polynomials and rational functions
We begin by recalling an elaboration of the Markov brothers inequality that is the basic principle used in the polynomial method [CGVTvH24]. The following convenient formulation may be found in [MdlS24b, Lem. 4.2].
Lemma 2.1.
For every real polynomial $P$ of degree at most $q$ and every $k\in\mathbf{Z}_{+}$,
$\sup_{t\in\left[0,\frac{1}{2q^{2}}\right]}|P^{(k)}(t)|\leq\frac{2^{2k+1}q^{4k}}{(2k-1)!!}\sup_{n\geq q^{2}}\left|P\left(\frac{1}{n}\right)\right|.$
We will require a variant of this principle for rational functions.
Lemma 2.2.
Let $P,Q$ be real polynomials of degree at most $q\in\mathbf{N}$, and let
$\Phi(t)=\frac{P(t)}{Q(t)}.$
Assume that there is a constant $C>1$ so that
$\bigg{|}P\bigg{(}\frac{1}{n}\bigg{)}\bigg{|}\leq C,\qquad\qquad C^{-1}\leq Q\bigg{(}\frac{1}{n}\bigg{)}\leq C$
for all $n\in\mathbf{N}$ with $n\geq(Cq)^{C}$. Then there is $C^{\prime}=C^{\prime}(C)>1$ so that for all $k\leq Cq$
$\sup_{t\in[0,(C^{\prime}q)^{-C^{\prime}}]}|\Phi^{(k)}(t)|\leq(C^{\prime}q)^{C^{\prime}k}.$
Proof.
Applying Lemma 2.1 with $q\leftarrow(Cq)^{C}$ yields
$\sup_{t\in[0,(2Cq)^{-2C}]}|P^{(k)}(t)|\leq 2C(2Cq)^{4Ck},\qquad\sup_{t\in[0,(2Cq)^{-2C}]}|Q^{(k)}(t)|\leq 2C(2Cq)^{4Ck},$
for all $k\in\mathbf{N}$. On the other hand, every $t\in[0,(2Cq)^{-2C-1}]$ is within distance at most $(2Cq)^{-4C-2}$ from some point $\frac{1}{n}$ with $n\geq(Cq)^{C}$. Thus
$\inf_{t\in[0,(2Cq)^{-2C-1}]}Q(t)\geq C^{-1}-(2Cq)^{-4C-2}\sup_{t\in[0,(2Cq)^{-2C-1}]}|Q^{(1)}(t)|\geq(2C)^{-1}.$
The conclusion now follows from the chain rule. Indeed, we can first estimate
$\sup_{t\in[0,(2Cq)^{-2C-1}]}\left|\left(\frac{1}{Q}\right)^{(k)}(t)\right|\leq(2C)^{2k+1}(2Cq)^{4Ck}k!$
as in the last part of [MdlS24b, proof of Lemma 7.2]. Thus
$|\Phi^{(k)}(t)|=\left|\sum_{i=0}^{k}\binom{k}{i}P^{(k-i)}(t)\left(\frac{1}{Q}\right)^{(i)}(t)\right|\leq 2^{k}(2C)^{2k+2}(2Cq)^{4Ck}k!$
for $t\in[0,(2Cq)^{-2C-1}]$. The conclusion follows as $k!\leq(Cq)^{k}$ for $k\leq Cq$. ∎
We finally recall a simple fact about trigonometric polynomials.
Lemma 2.3.
Let $f(\theta)=\sum_{k=0}^{q}a_{k}\cos(k\theta)$ and $m\in\mathbf{N}$. Then
$\sum_{k=1}^{q}k^{m-1}|a_{k}|\leq 4\|f^{(m)}\|_{[0,2\pi]}.$
Proof.
As $(k^{m}a_{k})_{0\leq k\leq q}$ are the Fourier coefficients of $f^{(m)}$, we have
$\sum_{k=1}^{q}k^{m-1}|a_{k}|\leq\left(\sum_{k=1}^{q}\frac{1}{k^{2}}\right)^{1/2}\left(\sum_{k=1}^{q}|k^{m}a_{k}|^{2}\right)^{1/2}\leq\sqrt{\frac{\pi^{2}}{6}}\|f^{(m)}\|_{L^{2}[0,2\pi]}$
by Cauchy-Schwarz and Parseval. We conclude using $\|f\|_{L^{2}[0,2\pi]}\leq\sqrt{2\pi}\|f\|_{[0,2\pi]}$. ∎
3 The generalized polynomial method
The aim of this section is to prove Theorem 1.5. The reader who is new to the polynomial method is encouraged to review [CGVTvH24, §2] for an introduction, before proceeding to the details of the proof.
In this section, we always assume that Assumptions 1.3 and 1.4 hold. We denote by $C$ the constant in Assumption 1.3, and by $r=|S|$ the size of the generating set of $\Lambda$.
The key step in the proof is to establish the following master inequality.
Lemma 3.1.
Fix a self-adjoint $x\in\mathbf{C}[\Lambda]$ and let $K=\|x\|_{C^{*}(\Lambda)}$. Then
$|u_{1}(h(x))|\leq r(6Cq|x|)^{6C}\|h\|_{[-K,K]}$
and
$\bigg{|}\mathbb{E}\big{[}{\mathop{\mathrm{tr}}h(\pi_{n}(x))}\big{]}-\tau\big{(}h(\lambda(x))\big{)}-\frac{1}{n}u_{1}(h(x))\bigg{|}\leq\frac{2r^{2}(6Cq|x|)^{12C}}{n^{2}}\|h\|_{[-K,K]}$
for all $n\geq 1$ and every real polynomial $h$ of degree at most $q$.
Proof.
We assume without loss of generality that $|x|\geq 1$ and $q\geq 1$ (otherwise $h(x)$ is a multiple of the identity, so both sides of the inequalities vanish). Let us write
$h(x)=\sum_{\gamma\in\Lambda}\alpha_{\gamma}\gamma.$
As $|h(x)|\leq q|x|$, there are at most $(2r+1)^{q|x|}\leq(3r)^{q|x|}$ nonzero coefficients $\alpha_{g}$. Applying Assumption 1.3 separately to each element of the support of $h(x)$ yields
$\Bigg{|}\mathbb{E}\big{[}{\mathop{\mathrm{tr}}h(\pi_{n}(x))}\big{]}-f_{h}\bigg{(}\frac{1}{n}\bigg{)}\Bigg{|}\leq\frac{(3Cq|x|)^{3Cq|x|}}{n^{q|x|+2}}\sum_{\gamma\in\Lambda}|\alpha_{\gamma}|$
for all $n\geq C(3q|x|)^{C}$, where we define the polynomial
$f_{h}(t)=\sum_{k=0}^{q|x|+1}u_{k}(h(x))\,t^{k}$
and we used that $q|x|+2\leq 3q|x|$. As
$\sum_{\gamma}|\alpha_{\gamma}|\leq(3r)^{q|x|/2}\|h(\lambda(x))\|_{L^{2}(\tau)}\leq(3r)^{q|x|/2}\|h\|_{[-K,K]}$ (3.1)
by Cauchy-Schwarz and as $\|\lambda(x)\|\leq K$, we readily estimate
$\Bigg{|}\mathbb{E}\big{[}{\mathop{\mathrm{tr}}h(\pi_{n}(x))}\big{]}-f_{h}\bigg{(}\frac{1}{n}\bigg{)}\Bigg{|}\leq\frac{1}{n^{2}}\|h\|_{[-K,K]}$ (3.2)
for $n\geq(3r)^{1/2}(3Cq|x|)^{3C}$. Now note that (3.2) implies the a priori bound
$\bigg{|}f_{h}\bigg{(}\frac{1}{n}\bigg{)}\bigg{|}\leq\big{|}\mathbb{E}\big{[}{\mathop{\mathrm{tr}}h(\pi_{n}(x))}\big{]}\big{|}+\frac{1}{n^{2}}\|h\|_{[-K,K]}\leq 2\|h\|_{[-K,K]}.$
Moreover, $u_{k}(h(x))$ are real as $h(x)$ is self-adjoint, so $f_{h}$ is a real polynomial. We can therefore apply Lemma 2.1 with $q^{2}\leftarrow(3r)^{1/2}(3Cq|x|)^{3C}$ to estimate
$\|f_{h}^{\prime}\|_{[0,r^{-1/2}(6Cq|x|)^{-3C}]}\leq r(6Cq|x|)^{6C}\|h\|_{[-K,K]},$
$\|f_{h}^{\prime\prime}\|_{[0,r^{-1/2}(6Cq|x|)^{-3C}]}\leq r^{2}(6Cq|x|)^{12C}\|h\|_{[-K,K]}.$
We can now readily conclude the proof. Indeed, as $u_{1}(h(x))=f_{h}^{\prime}(0)$, the first part of the statement is immediate. On the other hand, Taylor expanding $f_{h}$ yields
$\bigg{|}f_{h}\bigg{(}\frac{1}{n}\bigg{)}-u_{0}(h(x))-\frac{1}{n}u_{1}(h(x))\bigg{|}\leq\frac{1}{n^{2}}\|f_{h}^{\prime\prime}\|_{[0,\frac{1}{n}]}\leq\frac{r^{2}(6Cq|x|)^{12C}}{n^{2}}\|h\|_{[-K,K]}$
for $n\geq r^{1/2}(6Cq|x|)^{3C}$, so the second part follows in this case using (3.2) and that $u_{0}(x)=\tau(\lambda(x))$ by Assumption 1.3. Finally, when $n<r^{1/2}(6Cq|x|)^{3C}$, we have
$\bigg{|}\mathbb{E}\big{[}{\mathop{\mathrm{tr}}h(\pi_{n}(x))}\big{]}-\tau\big{(}h(\lambda(x))\big{)}-\frac{1}{n}u_{1}(h(x))\bigg{|}$ $\leq 2\|h\|_{[-K,K]}+\frac{1}{n}|u_{1}(h(x))|$
$\leq\frac{2r^{2}(6Cq|x|)^{12C}}{n^{2}}\|h\|_{[-K,K]}$
using the triangle inequality, the bound on $|u_{1}(h(x))|$, and $1\leq\frac{r^{1/2}(6Cq|x|)^{3C}}{n}$. ∎
With the statement of Lemma 3.1 in hand, the remainder of the argument is now precisely as in [CGVTvH24]. We first extend Lemma 3.1 to smooth $h$.
Corollary 3.2.
Fix a self-adjoint $x\in\mathbf{C}[\Lambda]$ and let $K=\|x\|_{C^{*}(\Lambda)}$. Then $\nu(h)\stackrel{{\scriptstyle\text{def}}}{{=}}u_{1}(h(x))$ extends to a compactly supported distribution, and
$\bigg{|}\mathbb{E}\big{[}{\mathop{\mathrm{tr}}h(\pi_{n}(x))}\big{]}-\tau\big{(}h(\lambda(x))\big{)}-\frac{1}{n}\nu(h)\bigg{|}\leq\frac{8r^{2}(6C|x|)^{12C}}{n^{2}}\|f^{(m)}\|_{[0,2\pi]}$
for all $n\geq 1$ and $h\in C^{\infty}(\mathbf{R})$. Here $m=1+\lceil 12C\rceil$ and $f(\theta)=h(K\cos\theta)$.
Proof.
That $\nu$ extends to a compactly supported distribution follows directly from the first part of Lemma 3.1 and [CGVTvH24, Lem. 4.7].
Now assume first that $h$ is a polynomial of degree $q$. Then we can uniquely write
$h(t)=\sum_{k=0}^{q}a_{k}T_{k}\bigg{(}\frac{t}{K}\bigg{)},$
where $T_{k}$ is the Chebyshev polynomial defined by $T_{k}(\cos(\theta))=\cos(k\theta)$. Applying the second part of Lemma 3.1 separately to each Chebyshev polynomial yields
$\bigg{|}\mathbb{E}\big{[}{\mathop{\mathrm{tr}}h(\pi_{n}(x))}\big{]}-\tau\big{(}h(\lambda(x))\big{)}-\frac{1}{n}u_{1}(h(x))\bigg{|}\leq\frac{2r^{2}(6C|x|)^{12C}}{n^{2}}\sum_{k=1}^{q}k^{12C}|a_{k}|.$
The conclusion now follows from Lemma 2.3 in the case that $h$ is a polynomial. Since the resulting bound does not depend on the degree of $h$ and as polynomials are dense in $C^{\infty}(\mathbf{R})$, the conclusion extends by continuity to any $h\in C^{\infty}(\mathbf{R})$. ∎
We can now state a quantitative version of Theorem 1.5.
Theorem 3.3.
Fix a self-adjoint $x\in\mathbf{C}[\Lambda]$ and let $K=\|x\|_{C^{*}(\Lambda)}$. Then
$\mathbb{P}\big{[}\|\pi_{n}(x)\|\geq\|\lambda(x)\|+\varepsilon\big{]}\leq\frac{r^{2}}{n}\bigg{(}\frac{c|x|K}{\varepsilon}\bigg{)}^{c}$
for every $n\geq 1$ and $\varepsilon>0$. Here $c$ is a constant that depends only on $C$.
Proof.
By [CGVTvH24, Lem. 4.10], there exists $h\in C^{\infty}(\mathbf{R})$ with values in $[0,1]$ so that
  1. 1.
    $h(z)=0$ for $|z|\leq\|\lambda(x)\|+\frac{\varepsilon}{2}$ and $h(z)=1$ for $|z|\geq\|\lambda(x)\|+\varepsilon$;
  2. 2.
    $\|f^{(m)}\|_{[0,2\pi]}\leq(\frac{cK}{\varepsilon})^{c}$ for $f(\theta)=h(K\cos\theta)$.
Moreover, Assumption 1.4 and [CGVTvH24, Lem. 4.9] yield $\mathop{\text{supp}}\nu\subseteq[-\|\lambda(x)\|,\|\lambda(x)\|]$, where $\nu$ is defined in Corollary 3.2. Thus $\tau(h(\lambda(x)))=\nu(h)=0$, so that
$\mathbb{P}\big{[}\|\pi_{n}(x)\|\geq\|\lambda(x)\|+\varepsilon\big{]}\leq\mathbb{E}\big{[}{\mathop{\mathrm{Tr}}h(\pi_{n}(x))}\big{]}\leq\frac{8r^{2}(6C|x|)^{12C}}{n}\bigg{(}\frac{cK}{\varepsilon}\bigg{)}^{c}$
using Corollary 3.2. This concludes the proof. ∎
The statement of Theorem 1.5 now follows immediately from Theorem 3.3 (if $x$ is not self-adjoint, we simply apply Theorem 3.3 to $x^{*}x$). However, Theorem 3.3 provides much stronger information, since it even yields a polynomial rate
$\|\pi_{n}(x)\|\leq\|\lambda(x)\|+O_{{\rm P}}(n^{-1/c}).$
Remarks
  1. 1.
    Assumption 1.3 ensures that $\mathbb{E}[\mathop{\mathrm{tr}}h(\pi_{n}(x))]$ can be approximated by a polynomial of $\frac{1}{n}$. At first sight, however, it is somewhat surprising that this suffices for the polynomial method. As is explained in [CGVTvH24, §2.3], the key feature of this method is that it captures certain delicate cancellations between the coefficients of $h(x)=\sum_{\gamma}\alpha_{\gamma}\gamma$. On the other hand, Assumption 1.3 only applies to each $\gamma$ separately and is therefore unable to preserve cancellations.
    The reason this does not matter is that the price for disregarding cancellations is exponential in $q|x|$ (see (3.1)), and can therefore be absorbed by the error term in Assumption 1.3 by expanding to order $\sim q|x|$. We can therefore first approximate the spectral statistics by high degree polynomials in the most naive manner, and then proceed to capture the cancellations in the latter using the polynomial method. This is crucial for applications, since we do not know how to establish a priori bounds as in Assumption 1.3 that capture cancellations.
  2. 2.
    While we developed the argument above in the group setting, the method is very robust and applies to many other situations. The only part of the argument that relied on the group structure is (3.1). In other situations, one can achieve a similar conclusion by applying a classical result of V. Markov which states that
    $\sum_{k=0}^{q}|a_{k}|\leq e^{q/K}\|h\|_{[-K,K]}$
    for every real polynomial $h(t)=\sum_{k=0}^{q}a_{k}t^{k}$ and $K>0$ [Tim63, §2.6, Eq. (9)], at the expense of a less explicit dependence of the constants on the choice of $x$.
4 Effective approximation by polynomials
The aim of this section is to prove Theorem 1.7.
Throughout this section, $C>0$ will denote a constant that may depend on the genus $g$, which can change from line to line but will only increase finitely many times.
4.1 Background
Representation theory of the symmetric group
A (number [^7] As opposed to a set partition.) partition of $n$ is a non-increasing sequence of non-negative integers
$\lambda=(\lambda_{1},\lambda_{2},\ldots,\lambda_{\ell(\lambda)})$ with
$\sum_{i=1}^{\ell(\lambda)}\lambda_{i}=n.$
In this case we write $\lambda\vdash n$ and say $\lambda$ has size $n$, $|\lambda|=n$. We think of partitions and Young diagrams (YDs) interchangeably in the sequel. If $\lambda$ is a YD we write $\lambda^{\vee}$ for the conjugate YD obtained by interchanging rows and columns. We write $b_{\lambda}$ for the number of boxes outside the first row of $\lambda$ and $b_{\lambda}^{\vee}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}b_{\lambda^{\vee}}$ the number of boxes outside the first column.
The equivalence classes of irreducible representations (irreps) of $S_{n}$ are parameterized (using Young symmetrizers, for example) by the YDs $\lambda\vdash n$. Given $\lambda\vdash n,$ we write $\chi_{\lambda}$ for the character of the associated irrep and $d_{\lambda}=\chi_{\lambda}(1)$ for its dimension. We write $V^{\lambda}$ for the corresponding $S_{n}$-module. As $\chi_{\lambda^{\vee}}=\mathrm{sign}\cdot\chi_{\lambda}$ we have $d_{\lambda}=d_{\lambda^{\vee}}$.
If one YD $\mu$ is contained in another $\lambda$, we say $\mu\subset\lambda$, and if they differ by $k$ boxes, we write $\mu\subset_{k}\lambda$. For YDs $\mu\subset_{k}\lambda$, a (standard) tableau of shape $\lambda/\mu$ is a filling of the $k$ boxes of $\lambda$ not in $\mu$ by the numbers in $[k]$ such that the numbers in each row (resp. column) are increasing from left to right (resp. top to bottom). We write $\mathrm{Tab}(\lambda/\mu)$ for the collection of such tableaux. It is obvious that when $\mu\subset_{k}\lambda$,
$|\mathrm{Tab}(\lambda/\mu)|\leq k!.$
We use the fact that every irrep of $S_{n}$ can be realized over the reals and therefore as orthogonal matrices (so no complex conjugates appear in our calculations).
The Witten zeta function
The Witten zeta function of $S_{n}$ is defined by
$\zeta(s;S_{n})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{\lambda\vdash n}\frac{1}{d_{\lambda}^{s}}.$
This appears as a normalization factor in the uniform measure on $\mathrm{Hom}(\Gamma_{g},S_{n})$ and hence is important in the sequel. By a result of Liebeck and Shalev [LS04, Prop. 2.5] (independently, Gamburd [Gam06, Prop. 4.2]), for real $s>0$,
$\zeta(s;S_{n})=2+O_{s}\left(n^{-s}\right)$ (4.1)
as $n\to\infty$. We also have the following effective tail bound [MNP22, Prop. 4.6].
Proposition 4.1.
For all $s>0$, there is $\kappa=\kappa(s)>1$ such that
$\sum_{\begin{subarray}{c}\lambda\vdash n\\ b_{\lambda},b_{\lambda}^{\vee}\geq b\end{subarray}}\frac{1}{d_{\lambda}^{s}}\leq\left(\frac{\kappa b^{2s}}{(n-b^{2})^{s}}\right)^{b}$ (4.2)
for all $b,n\in\mathbf{N}$ with $b^{2}\leq\frac{n}{3}$.
Expansion of $\mathbb{E}_{g,n}[\mathsf{fix}_{\gamma}]$ by expected numbers of embeddings
In the following, $\phi$ is a uniformly random element of $\mathrm{Hom}(\Gamma_{g},S_{n})$. We now recap on the language of [MP23]. This is just a formalism for expressing the facts:
  • $\mathbb{E}_{g,n}[\mathsf{fix}_{\gamma}]$ is $n$ times the probability that $\phi(\gamma)$ fixes a given point, say $1\in[n]$, by invariance under $S_{n}$ conjugation.
  • The event $\mathcal{E}_{\gamma}$ that $\phi(\gamma)$ fixes $1$ can be partitioned into sub-events in a variety of different ways.
In [MP23] these sub-events are chosen more carefully than we have to here.
We use the standard set of generators $a_{1},\ldots,b_{g}$ of $\Gamma_{g}$ from (1.2). The partitioning is described by reference to the Schreier graph
$X_{\phi}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathrm{Schreier}\left(\{\phi(a_{1}),\ldots,\phi(b_{g})\},[n]\right)$
defined by the action of $\phi(a_{1}),\ldots,\phi(b_{g})$ on $[n]$, i.e., there is a directed edge from $i$ to $j$ for each $f\in\{a_{1},\ldots,b_{g}\}$ such that
$\phi(f)[i]=j.$
Every directed edge is colored by the generator from which it arises. [^8] The definition of $X_{\phi}$ given here is the 1-skeleton of the $X_{\phi}$ of [MP23], which includes $n$ additional 2-cells to form a closed surface.
Given $\gamma\in\Gamma_{g}$, let $\gamma^{\prime}$ be a conjugate of $\gamma$ whose shortest representing word in the generators is cyclically reduced. We have $|\gamma^{\prime}|\leq|\gamma|$ and
$\mathbb{E}_{g,n}[\mathsf{fix}_{\gamma}]=\mathbb{E}_{g,n}[\mathsf{fix}_{\gamma^{\prime}}],$
so without loss of generality in the following suppose $\gamma$ is cyclically reduced.
Let $F_{2g}$ denote the free group on $a_{1},\ldots,b_{g}$. Let $C_{\gamma}$ denote a circle subdivided into $|\gamma|$ intervals (viewed as an undirected graph). Direct and label these intervals by $a_{1},\ldots,b_{g}$ according to an expression of $\gamma$ of length $|\gamma|$ in these generators — the direction indicates whether a generator or its inverse appears at a given location.
Let $\mathcal{R}$ denote the collection of surjective labeled-graph morphisms
$r:C_{\gamma}\twoheadrightarrow W_{r}$
such that:
  1. 1.
    $W_{r}$ is folded in the sense that every vertex has at most one incoming $f$-labeled half-edge and at most one outgoing $f$-labeled half-edge, for each $f\in\{a_{1},\ldots,b_{g}\}$. (Note that the assumption $\gamma$ is cyclically reduced means $C_{\gamma}$ itself is folded.)
  2. 2.
    Every path in $W_{r}$ spelling out an element of $F_{2g}$ that is in the kernel of $F_{2g}\to\Gamma_{g}$ is closed.
In the language of [MP22, MP23] these $W_{r}$ are (very special cases of) tiled surfaces. [^9] In (ibid.) it is a canonical thickening of the graph that is thought of as the surface.
Any homomorphism of folded labeled graphs $C_{\gamma}\to X_{\phi}$ factors uniquely as a surjective morphism followed by an injective one, namely
$C_{\gamma}\stackrel{{\scriptstyle r}}{{\twoheadrightarrow}}W_{r}\hookrightarrow X_{\phi},$
for unique $r\in\mathcal{R}$. In the language of [MP23], this factoring property means $\mathcal{R}$ is a resolution of $C_{\gamma}$.
Remark 4.2 (Comparison to [MP23]).
Although the resolution we define here might be the most natural, the methods of [MP23] leading to [MP23, Thm. 1.2] — not proved here — only work on $W_{r}$ who have, very roughly speaking, geodesic boundary. So in [MP23] great care is taken over finding a finer resolution whose boundary components have desired properties. For our purposes, the simplest resolution suffices.
We have
$|\mathcal{R}|\leq|\gamma|!$ (4.3)
since each element is defined by the partition of the vertices of $C_{\gamma}$ given by the fibers of the surjective immersion, and there are $|\gamma|$ vertices. By [MP23, Lem. 2.7 and 2.9]
$\mathbb{E}_{g,n}[\mathsf{fix}_{\gamma}]=\sum_{r\in\mathcal{R}}\mathbb{E}_{n}^{\mathrm{\mathrm{emb}}}(W_{r}),$ (4.4)
where
$\mathbb{E}_{n}^{\mathrm{emb}}(W_{r})=\mathbb{E}_{\phi\in\mathbb{X}_{g,n}}\{\,\#\text{ injective morphisms $W_{r}\to X_{\phi}\,\}.$}$
In the sequel, for $W_{r}$ as above we write $\mathfrak{v}(W_{r}),\mathfrak{e}(W_{r})$ for the number of vertices and edges of $W_{r}$. For $f\in\{a_{1},\ldots,b_{g}\}$ we write $\mathfrak{e}_{f}(W_{r})$ for the number of $f$-labeled edges of $W_{r}$. With things as before, clearly
$\mathfrak{v}(W_{r}),\mathfrak{e}(W_{r}),\mathfrak{e}_{f}(W_{r})\leq|\gamma|.$
Expected number of embeddings, prior results
In the remainder of §4, we restrict to $g=2$ for simplicity of exposition. We now use $a,b,c,d$ for $a_{1},b_{1},a_{2},b_{2}$. The only things that depend on $g$ in a non-obvious way are constants and we indicate how these depend on $g$ throughout. All integrals in the rest of this section are taken with respect to uniform probability measures.
Let $Y$ denote some fixed $W_{r}$ from our resolution $\mathcal{R}$. Let $\mathfrak{v}=\mathfrak{v}(Y)$, $\mathfrak{e}_{f}=\mathfrak{e}_{f}(Y)$. In the following we label the vertices of $Y$ injectively by $1,\ldots,\mathfrak{v}$.
For $g_{f}\in S_{n}$, $f\in\{a,b,c,d\}$, we say $g_{f}$ obey $Y$ if the fixed labeling of the vertices of $Y$ induces an embedding
$Y\hookrightarrow\mathrm{Schreier}\left(\{g_{f}\},[n]\right),$
or in simple terms, if $Y$ has an $f$-colored directed edge from a vertex labeled $i$ to vertex labeled $j$, then $g_{f}(i)=j$.
We have by [MP23, (2.1), (5.6), and proof of Prop. 5.1], since $\mathfrak{v}(Y)\leq q$, for $n\geq q$
$\mathbb{E}_{n}^{\mathrm{emb}}(Y)=\frac{1}{\zeta(2;S_{n})}\frac{(n)_{\mathfrak{v}}}{\prod_{f\in a,b,c,d}(n)_{\mathfrak{e}_{f}}}\sum_{\lambda\vdash n}d_{\lambda}\Theta_{\lambda}\left(Y\right),$
where
$\Theta_{\lambda}\left(Y\right)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\left(\prod_{f\in a,b,c,d}(n)_{\mathfrak{e}_{f}}\right)\int_{g_{f}\in S_{n}}\mathbf{1}\{g_{f}\text{ obey $Y$\}}\chi_{\lambda}\left([g_{a},g_{b}][g_{c},g_{d}]\right)$ (4.5)
(there is a change of normalization since we pass between integrals over different groups: in the notation of (ibid.) $|G_{f}|=(n-\mathfrak{e}_{f})!$).
Remark 4.3.
In [MP23] the definition of $\Theta_{\lambda}\left(Y\right)$ depends on an auxiliary labeling $\mathcal{J}_{n}:Y^{(0)}\to\{n-\mathfrak{v}+1,\ldots,n\}$. This is not needed here, as the definition of $\Theta_{\lambda}\left(Y\right)$ in (4.5) is the same for any labeling by invariance of the uniform measure on $S_{n}$ by conjugation.
From [MP23, Prop. 5.8]
$\Theta_{\lambda}(Y)=$ $\sum_{\nu\subset_{\mathfrak{v}}\lambda}d_{\nu}\sum_{\begin{subarray}{c}\nu\subset\mu_{f}\subset_{\mathfrak{e}_{f}}\lambda\\ f\in\{a,b,c,d\}\end{subarray}}\frac{1}{d_{\mu_{a}}d_{\mu_{b}}d_{\mu_{c}}d_{\mu_{d}}}\Upsilon_{n}\left(\left\{\sigma_{f}^{\pm},\tau_{f}^{\pm}\right\},\nu,\left\{\mu_{f}\right\},\lambda\right),$ (4.6)
where
$\Upsilon_{n}\left(\left\{\sigma_{f}^{\pm},\tau_{f}^{\pm}\right\},\nu,\left\{\mu_{f}\right\},\lambda\right)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{\begin{gathered}r_{f}^{+},r_{f}^{-}\in\mathrm{Tab}\left(\mu_{f}/\nu\right)\\ s_{f},t_{f}\in\mathrm{Tab}\left(\lambda/\mu_{f}\right)\end{gathered}}\mathcal{M}\left(\left\{\sigma_{f}^{\pm},\tau_{f}^{\pm},r_{f}^{\pm},s_{f},t_{f}\right\}\right)$ (4.9)
and ($\mathcal{M}\left(\left\{\sigma_{f}^{\pm},\tau_{f}^{\pm},r_{f}^{\pm},s_{f},t_{f}\right\}\right)$ is a product of matrix coefficients of unitary operators on unit vectors as in [MP23, eq. (5.15)]. All we need here about this product is that
|($\left|\mathcal{M}\left(\left\{\sigma_{f}^{\pm},\tau_{f}^{\pm},r_{f}^{\pm},s_{f},t_{f}\right\}\right)\right|\leq 1.$ (4.10)
4.2 Tail estimate for $\mathbb{E}_{n}^{\mathrm{\mathrm{emb}}}(W_{r})$
In this section we prove:
Proposition 4.4.
There is $C>0$ such that if $\mathfrak{v}(Y)\leq q$ and $n\geq 28q^{2}$ then
$\mathbb{E}_{n}^{\mathrm{emb}}(Y)$ $=\frac{1}{\zeta(2;S_{n})}\frac{(n)_{\mathfrak{v}}}{\prod_{f\in a,b,c,d}(n)_{\mathfrak{e}_{f}}}\sum_{\begin{subarray}{c}\lambda\vdash n\\ \min(b_{\lambda},b_{\lambda}^{\vee})\leq 4q\end{subarray}}d_{\lambda}\Theta_{\lambda}\left(Y\right)$
$\,\,+O\left((Cq)^{Cq}n^{-q}\right).$
Proof.
Using $\left|\mathrm{Tab}\left(\mu_{f}/\nu\right)\right|,\left|\mathrm{Tab}\left(\lambda/\mu_{f}\right)\right|\leq\mathfrak{v}!\leq q!$ together with (4.10) in (4.6) we obtain
$\Theta_{\lambda}(Y)\leq(Cq)^{Cq}\sum_{\nu\subset_{\mathfrak{v}}\lambda}\frac{1}{d_{\nu}^{3}}.$
Note if $b_{\lambda}>4q$ then $b_{\nu}\geq b_{\lambda}-\mathfrak{v}>3q$. We have then
$\sum_{\begin{subarray}{c}\lambda\vdash n\\ b_{\lambda},b_{\lambda}^{\vee}>4q\end{subarray}}d_{\lambda}\Theta_{\lambda}\left(Y\right)\leq(Cq)^{Cq}\sum_{\lambda\vdash n}\sum_{\begin{subarray}{c}\nu\subset_{\mathfrak{v}}\lambda\\ b_{\nu},b_{\nu}^{\vee}>3q\end{subarray}}\frac{d_{\lambda}}{d_{\nu}^{3}}.$
As in [MP23, proof of Lemma 5.23] the above is
$\leq(n)_{\mathfrak{v}}(Cq)^{Cq}\sum_{\begin{subarray}{c}\nu\vdash n-\mathfrak{v}\\ b_{\nu},b_{\nu}^{\vee}>3q\end{subarray}}\frac{1}{d_{\nu}^{2}}.$ (4.11)
Hence by (4.2)
$\sum_{\begin{subarray}{c}\nu\vdash n-\mathfrak{v}\\ b_{\nu},b_{\nu}^{\vee}>3q\end{subarray}}\frac{1}{d_{\nu}^{2}}\leq C\left(\frac{(Cq)^{4}}{(n-\mathfrak{v}-9q^{2})^{2}}\right)^{3q}\leq C(Cq)^{Cq}n^{-6q}$ (4.12)
in e.g. $n\geq 28q^{2}$. (Note that to use (4.2) for $b=3q$ as above we need $\mathfrak{v}+3(3q)^{2}\leq n$.) Combining all previous arguments gives
$\mathbb{E}_{n}^{\mathrm{emb}}(Y)$ $=\frac{1}{\zeta(2;S_{n})}\frac{(n)_{\mathfrak{v}}}{\prod_{f\in a,b,c,d}(n)_{\mathfrak{e}_{f}}}\sum_{\begin{subarray}{c}\lambda\vdash n\\ \min(b_{\lambda},b_{\lambda}^{\vee})\leq 4q\end{subarray}}d_{\lambda}\Theta_{\lambda}\left(Y\right)$
$+O\left(\frac{1}{\zeta(2;S_{n})}\frac{(n)_{\mathfrak{v}}^{2}}{\prod_{f\in a,b,c,d}(n)_{\mathfrak{e}_{f}}}(Cq)^{Cq}n^{-6q}\right).$
The denominator Pochhammer symbols in the error are $\geq 1$ and numerators $\leq n^{2q}$. Hence (also using (4.1)) the whole error is on the order of
$(Cq)^{Cq}n^{-q}$
as required.
(For general $g\geq 2$, the number of tableaux being summed over in the combination of (4.6) and (4.9) is $8g$, hence $C$ depends linearly on $g$, the summand of (4.11) is $\frac{1}{d_{\nu}^{2g-2}}\leq\frac{1}{d_{\nu}^{2}}$ and the estimate (4.12) can be kept the same. The rest of the proof is the same, changing the product $\prod_{f\in a,b,c,d}(n)_{\mathfrak{e}_{f}}$ to a product over $2g$ generators and $\zeta(2;S_{n})$ to $\zeta(2g-2;S_{n}$).) ∎
4.3 Rationality of contributions from fixed representations
Given a YD $\lambda\vdash b$, for $n\geq 2b$ we write $\lambda^{+}(n)$ for the YD of size $n$ with $\lambda$ outside the first row. Our aim here is to prove the following rationality result.
Proposition 4.5.
Let $b\in\mathbf{Z}_{+}$, $n\geq\mathfrak{v}(Y)+9b$, and $\lambda\vdash b$. As a function of $n$, $\Theta_{\lambda^{+}(n)}(Y)$ agrees with a rational function of $n$ with coefficients in $\mathbf{Q}$ whose denominator can be taken to be
$(n)_{\mathfrak{v}+2b}^{5}.$
The proof of Proposition 4.5 uses a result of Cassidy [Cas23]. Let $\{e_{i}\}_{i=1}^{n}$ denote the standard orthonormal basis of $\mathbf{C}^{n}$ with its standard inner product. Given a set partition $\pi$ of $[2b]$, define the following endomorphism $P_{\pi}$ of $(\mathbf{C}^{n})^{\otimes b}$ by
$\langle P_{\pi}(e_{i_{1}}\otimes\cdots\otimes e_{i_{b}}),e_{i_{b+1}}\otimes\cdots\otimes e_{i_{2b}}\rangle\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\begin{cases}1&\text{if $i_{p}=i_{q}$ iff $p$ and $q$ in same block of $\pi$}\\ 0&\text{else.}\end{cases}$
Denote by $\mathrm{Part}([2b])$ the set partitions of $[2b]$, and by $\iota$ the inclusion $S_{b}\hookrightarrow\mathrm{Part}([2b])$ defined by viewing an element of $S_{b}$ as a matching between two horizontal rows of $b$ elements, the first row labeled by $[b]$ and the second row labeled by $[2b]\backslash[b]$. Say $\pi_{1}\leq\pi_{2}$ if the blocks of $\pi_{1}$ subdivide those of $\pi_{2}$. Let $\mathrm{SubPerm}(b)\subseteq\mathrm{Part}([2b])$ denote “sub-permutations”: set partitions $\pi$ of $[2b]$ such that $\pi\leq\iota(\sigma)$ for some $\sigma\in S_{b}$, namely, such that every block of $\pi$ contains at most one element from each of the two horizontal rows. We write $|\pi|$ for the number of blocks of $\pi$. Here is Cassidy’s formula. [^10] We point out that we do not use the full force of Cassidy’s work in this paper, and it can be bypassed by the (stable) character theory of $S_{n}$. In particular, all we need here is to know that for large $n$, $\chi_{\lambda^{+}(n)}(g)$ is a fixed polynomial of $\mathsf{fix}(g^{k})$ for finitely many $k$ (depending on $\lambda$) where the polynomial has rational coefficients of $n$ with denominators that are ‘not too bad’. The more classical formula [HP23, (B.1)] would accomplish the same thing. Using Cassidy’s result is handy here, however, and we believe that this formalism will also be useful for some future investigations in this area. (This differs from the formula in [Cas23] by Möbius inversion and orthogonality.)
Theorem 4.6 (Cassidy [Cas23, Thm. 1.2]).
For $b\in\mathbf{Z}_{+}$, $n\geq 2b$, and $\lambda\vdash b$,
$\mathfrak{p}_{\lambda}=(-1)^{b}d_{\lambda^{+}(n)}\sum_{\pi\in\mathrm{SubPerm}(b)}\frac{(-1)^{|\pi|}}{(n)_{|\pi|}}\left(\sum_{\begin{subarray}{c}\tau\in S_{b}\\ \pi\leq\iota(\tau)\end{subarray}}\chi_{\lambda}(\tau)\right)P_{\pi}$ (4.13)
is the orthogonal projection in $(\mathbf{C}^{n})^{\otimes b}$ onto $d_{\lambda}$ copies of $V^{\lambda^{+}(n)}$ as an $S_{n}$-representation.
Proof of Proposition 4.5.
Here we evaluate $\Theta_{\lambda^{+}(n)}\left(Y\right)$ differently to [MP23]. If
$\rho_{b}:S_{n}\to\mathrm{End}\left(\left(\mathbf{C}^{n}\right)^{\otimes b}\right)$
is the $b$th tensor power of the standard permutation representation of $S_{n}$, we can rewrite (4.5) as
$\Theta_{\lambda^{+}(n)}\left(Y\right)=\frac{\prod_{f\in a,b,c,d}(n)_{\mathfrak{e}_{f}}}{d_{\lambda}}\int_{g_{f}\in S_{n}}\mathbf{1}\{g_{f}\text{ obey \,$Y$)\}}{\mathop{\mathrm{Tr}}}_{(\mathbf{C}^{n})^{\otimes b}}\left(\mathfrak{p}_{\lambda}\rho_{b}\left([g_{a},g_{b}][g_{c},g_{d}]\right)\right)$
where $\mathfrak{p}_{\lambda}$ is the projection operator from Theorem 4.6.
From (4.13), $\mathfrak{p}_{\lambda}$ is a linear combination of partition operators $P_{\pi};$ all the coefficients are rational functions of $n$ in $n\geq 2b$ and can be put over a common denominator of $(n)_{2b}$. Here we used that $d_{\lambda^{+}(n)}$ is a polynomial of $n$ in this range by the hook-length formula [FRT54]. From this we learn $\Theta_{\lambda^{+}(n)}\left(Y\right)$ is a linear combination of
$\int_{g_{f}\in S_{n}}\mathbf{1}\{g_{f}\text{ obey \,$Y$\}}{\mathop{\mathrm{Tr}}}_{(\mathbf{C}^{n})^{\otimes b}}\left(P_{\pi}\rho_{b}\left([g_{a},g_{b}][g_{c},g_{d}]\right)\right)$ (4.14)
with rational coefficients in $n$, with common denominator $(n)_{2b}$.
We can now expand
${\mathop{\mathrm{Tr}}}_{(\mathbf{C}^{n})^{\otimes b}}\left(P_{\pi}\rho_{b}\left([g_{a},g_{b}][g_{c},g_{d}]\right)\right)=\\ \sum_{I,J,K,L,M,N,O,Q,R}(P_{\pi})_{RI}(g_{a})_{IJ}(g_{b})_{JK}(g_{a})_{LK}(g_{b})_{ML}(g_{c})_{MN}(g_{d})_{NO}(g_{c})_{QO}(g_{d})_{RQ}$ (4.15)
with $I,J,K,L,M,N,O,Q,R\in[n]^{b}$, where by slight abuse of notation we write $(g_{f})_{IJ}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\rho_{b}(g_{f})_{IJ}$. We view $I,J,K,L,M,N,O,Q,R$ as well as the fixed labeling of $Y^{(0)}$ concatenated together as a function
$\Omega:Y^{(0)}\cup[9b]\to[n].$
The function $\Omega$ and all these indices contain exactly the same information. We now partition the summation in (4.15) according to the partition $p_{\Omega}$ of $Y^{(0)}\cup[9b]$ that the function $\Omega$ defines: two locations are in the same block of $p_{\Omega}$ exactly when they have the same image. Write $\pi\to\Omega$ if and only if the values of $R,I$ encoded by $\Omega$ satisfy $(P_{\pi})_{RI}=1$. We obtain
$\mathbf{1}\{g_{f}\text{ obey \,$Y$\}}{\mathop{\mathrm{Tr}}}_{(\mathbf{C}^{n})^{\otimes b}}\left(P_{\pi}\rho_{b}\left([g_{a},g_{b}][g_{c},g_{d}]\right)\right)$
$=\sum_{\pi\to\Omega}\mathbf{1}\{g_{f}\text{ obey \,$Y$\}}(g_{a})_{IJ}(g_{b})_{JK}(g_{a})_{LK}(g_{b})_{ML}(g_{c})_{MN}(g_{d})_{NO}(g_{c})_{QO}(g_{d})_{RQ}.$ (4.16)
In the above sum the labeling of $Y^{(0)}$ is fixed as usual. Now we use the following elementary integration formula:
$\int_{g\in S_{n}}g_{i_{1}j_{1}}g_{i_{2}j_{2}}\cdots g_{i_{k}j_{k}}$
=$=\begin{cases}\frac{1}{(n)_{|\theta|}}&\text{if the maps $\ell\mapsto i_{\ell}$ and $\ell\mapsto j_{\ell}$ define the same partition $\theta\in\mathrm{Part}([k])$}\\ 0&\text{otherwise.}\end{cases}$
Notice that $\mathbf{1}\{g_{f}\text{ obey \,$Y$\}}$ is a product of matrix entries of the $g_{f}$:
$\mathbf{1}\{g_{f}\text{ obey \,$Y$\}}=\prod_{f\in\{a,b,c,d\}}\prod_{\{(i,j)\,:\,Y\text{ has $f$-colored edge from $i\to j$\}}}(g_{f})_{ij}.$
Hence for $n\geq\mathfrak{v}+9b$ the integral of the summand of (4.16) depends only on $p_{\Omega}$ and is either zero or equal to
$\frac{1}{\prod_{f\in\{a,b,c,d\}}(n)_{|(p_{\Omega})_{f}|}}$ (4.17)
where $(p_{\Omega})_{f}$ is the partition induced by $p_{\Omega}$ on
  • vertices of $Y$ with an outgoing $f$-colored edge, together with
  • elements of $[9b]\backslash Y^{(0)}$ who are in the domain of indices at the start of the matrix coefficient, e.g. for $f=a$, domain of $I$ or domain of $L$.
Note that each
$|(p_{\Omega})_{f}|\leq\mathfrak{v}+2b.$
Therefore the total contribution from a fixed $p_{\Omega}$ to (4.14) is (4.17) times the number of $\Omega$ inducing $p_{\Omega}$. If a block of $p_{\Omega}$ contains an element of $Y^{(0)}$ then the corresponding value of $\Omega$ is fixed by the fixed labeling of $Y^{(0)}$. The other blocks have values in $[n]$ that must be disjoint from each other, hence the number of $\Omega$ for a given $p_{\Omega}$ is
$(n)_{\text{num. blocks of $p_{\Omega}$ disjoint from $Y^{(0)}$}}.$
Putting the previous arguments together proves the result.
(For general $g$ the only change is the number of factors $(n)_{|(p_{\Omega})_{f}|}$ in the denominator and the end result is a denominator $(n)_{\mathfrak{v}+2b}^{1+2g}$.) ∎
4.4 Proof of Theorem 1.7
We can now complete the proof.
Proof of Theorem 1.7.
Suppose that $|\gamma|\leq q$. We write $\zeta(n)=\zeta(2;S_{n})$. Now by (4.3), (4.4), and Proposition 4.4, for $n\geq 28q^{2}$
$\mathbb{E}_{g,n}[\mathsf{fix}_{\gamma}]$ $=\frac{1}{\zeta(n)}\sum_{r\in\mathcal{R}}\frac{(n)_{\mathfrak{v}(W_{r})}}{\prod_{f\in a,b,c,d}(n)_{\mathfrak{e}_{f}(W_{r})}}\sum_{\begin{subarray}{c}\lambda\vdash n\\ \min(b_{\lambda},b_{\lambda}^{\vee})\leq 4q\end{subarray}}d_{\lambda}\Theta_{\lambda}\left(W_{r}\right)$
$\,+O\left((Cq)^{Cq}n^{-q}\right).$
Since above either $b_{\lambda}\leq 4q$ or $b_{\lambda}^{\vee}\leq 4q$, for $n\geq Cq$ these events cannot happen simultaneously. Moreover, it is easy to check from (4.5) that $d_{\lambda}\Theta_{\lambda}\left(W_{r}\right)$ is invariant under $\lambda\mapsto\lambda^{\vee}$ and hence our main term can be written
$\frac{2}{\zeta(n)}\sum_{r\in\mathcal{R}}\frac{(n)_{\mathfrak{v}(W_{r})}}{\prod_{f\in a,b,c,d}(n)_{\mathfrak{e}_{f}(W_{r})}}\sum_{|\lambda|\leq 4q}d_{\lambda^{+}(n)}\Theta_{\lambda^{+}(n)}\left(W_{r}\right).$
By the hook-length formula [FRT54] each $d_{\lambda^{+}(n)}$ is a divisor in $\mathbf{Q}[n]$ of
$(n)_{8q}=n(n-1)\cdots(n-8q+1)$
and in particular, a polynomial of $n$ in the range we consider. Since for everything above, $\mathfrak{v}(W_{r}),\mathfrak{e}_{f}(W_{r})\leq q,$ for $n\geq Cq$ Proposition 4.5 tells us that
$\Phi_{\gamma}(n)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{r\in\mathcal{R}}\frac{(n)_{\mathfrak{v}(W_{r})}}{\prod_{f\in a,b,c,d}(n)_{\mathfrak{e}_{f}(W_{r})}}\sum_{|\lambda|\leq 4q}d_{\lambda^{+}(n)}\Theta_{\lambda^{+}(n)}\left(W_{r}\right)$
is a rational function of $n$ whose denominator can be taken to be $(n)_{9q}^{9}.$ (For general $g$, this is replaced by $(n)_{9q}^{1+4g}$.)
So we write
$\Phi_{\gamma}(n)=\frac{p(n)}{(n)_{9q}^{9}}=\frac{1}{2}\zeta(n)\mathbb{E}_{g,n}[\mathsf{fix}_{\gamma}]+O\left((Cq)^{Cq}n^{-q}\right)$
using $\zeta\to 2$ as $n\to\infty$ (see (4.1)) to remove $\zeta$ from the error. Using the same fact again with a priori bound $|\mathbb{E}_{g,n}[\mathsf{fix}_{\gamma}]|\leq n$ we learn that $|\frac{1}{n}\frac{p(n)}{(n)_{9q}^{9}.}|$ is bounded as $n\to\infty$ and hence
$\deg(p)\leq 81q+1.$
Let
$g_{q}(t)=\prod_{k=0}^{9q-1}\left(1-kt\right)^{9}$
and for $t\stackrel{{\scriptstyle\mathrm{def}}}{{=}}n^{-1}$ write
$\Phi_{\gamma}(n)=\frac{p(n)}{(n)_{9q}^{9}}=\frac{n^{\deg(p)}}{n^{81q}}\frac{P(t)}{g_{q}(t)}=t^{81q-\deg(p)}\frac{P(t)}{g_{q}(t)}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{1}{t}\frac{Q(t)}{g_{q}(t)}.$
where $P(t)$ is a polynomial of $t$ of degree $\leq\deg(p)$. For $\tau\in[0,\frac{1}{Cq^{2}}]$ we have
$g_{q}(\tau)\geq C^{-1}$ (4.18)
so we have
$Q(t)=t^{81q-\deg(p)+1}P(t)=O(1)$ (4.19)
for $t\in\left[0,\frac{1}{Cq^{2}}\right]\cap\mathbf{N}^{-1}$ (using the a priori bound again). Note that $\deg(Q)\leq 81q+1$. (By the previous remarks, for general $g$, $C$ can be taken to be linear in $g$.)
To conclude, we begin with
$\mathbb{E}_{g,n}[\mathsf{fix}_{\gamma}]=\frac{2}{t\zeta(n)}\frac{Q(t)}{g_{q}(t)}+O\left((Cq)^{Cq}n^{-q}\right).$ (4.20)
We now need to deal with the rogue $\zeta(n)$ factor. We can use a sort of trick to do this efficiently. If $\gamma=\mathrm{id}$ then all the previous arguments apply to give
$t^{-1}=n=\mathbb{E}_{g,n}[\mathsf{fix}_{\mathrm{id}}]=\frac{2}{\zeta(n)}t^{-1}\frac{Q_{\mathrm{id}}(t)}{g_{q}(t)}+O\left((Cq)^{Cq}n^{-q}\right)$
where $Q_{\mathrm{id}}$ is a polynomial that also satisfies (4.19). Rearranging,
$\frac{Q_{\mathrm{id}}(t)}{g_{q}(t)}$ $=\zeta(n)\left(\frac{1}{2}+O\left((Cq)^{Cq}n^{-q-1}\right)\right).$ (4.21)
So when $n\geq(Cq)^{2C}$ we have
$\frac{2}{\zeta(n)}=\frac{g_{q}(t)}{Q_{\mathrm{id}}(t)}\left(1+O\left((Cq)^{Cq}n^{-q-1}\right)\right).$
Therefore
$\mathbb{E}_{g,n}[\mathsf{fix}_{\gamma}]=t^{-1}\frac{Q(t)}{Q_{\mathrm{id}}(t)}\left(1+O\left((Cq)^{Cq}n^{-q-1}\right)\right)+O\left((Cq)^{Cq}n^{-q}\right).$
Note that $Q_{\mathrm{id}}\geq C^{-1}$ in $n\geq q^{C}$ from (4.21) and (4.18). So by the bound on $Q$ from (4.19) (also for $Q_{\mathrm{id}}$) we obtain from Lemma 2.2 that for $\tau\in[0,\frac{1}{Cq^{2C}}]$
$\left|\left(\frac{Q}{Q_{\mathrm{id}}}\right)^{(i)}(\tau)\right|\leq C(Cq)^{Ci}.$
Therefore by Taylor’s theorem there are $a_{-1},a_{0},\ldots,a_{q}$ (depending on $\gamma)$ with
$\left(\frac{Q}{Q_{\mathrm{id}}}\right)(\tau)=a_{-1}+a_{0}\tau+\cdots+a_{q-1}\tau^{q}+O\left((Cq)^{Cq}\tau^{q+1}\right)$
for $\tau\in[0,\frac{1}{Cq^{2C}}]$. Thus for $n\geq Cq^{C}$
$\mathbb{E}_{g,n}[\mathsf{fix}_{\gamma}]=a_{-1}n+a_{0}+a_{1}n^{-1}+\cdots+a_{q-1}n^{-(q-1)}+O\left((Cq)^{Cq}n^{-q}\right).$
This concludes the proof. ∎
5 Geometry of proper powers
The aim of this section is to prove Theorem 1.8.
5.1 Embedding the Cayley graph of $\Gamma$ in $\mathbb{H}^{2}$
Denote by $\text{Cay}(\Gamma)$ the Cayley graph of $\Gamma$ with respect to the generators $\{a_{1},b_{1},\ldots,a_{g},b_{g}\}$ in (1.2). We use the following standard embedding of $\text{Cay}(\Gamma)$ in the hyperbolic plane $\mathbb{H}^{2}$, which is useful, inter alia, as it agrees with the description in [BS87].
$a_{1}$$b_{1}$$a_{1}$$b_{1}$$a_{2}$$b_{2}$$a_{2}$$b_{2}$
Figure 5.1: Illustration of the tiling $\mathcal{T}$ and embedded Cayley graph of $\Gamma$ for $g=2$.
Consider the tiling ${\cal T}$ of $\mathbb{H}^{2}$ by regular $4g$-gons with geodesic sides and interior angles $\frac{2\pi}{4g}$. The Cayley graph is then the tiling dual to ${\cal T}$. More precisely, the vertices of $\text{Cay}(\Gamma)$ are the centers of the $4g$-gons, and the edges are geodesic arcs between the centers of any two $4g$-gons sharing a side. Every edge of $\text{Cay}(\Gamma)$ is directed and labeled by one of the $2g$ generators. For any generator $x$, when traversing an $x$-edge against its direction, one reads $x^{-1}$. At each vertex, there is exactly one edge directed outward and one directed inward with every given label. The cyclic order of the $4g$ outgoing edges, say clockwise, is
$a_{1},b_{1}^{-1},a_{1}^{-1},b_{1},a_{2},b_{2}^{-1},a_{2}^{-1},b_{2},\ldots,b_{g}.$
(Up to isometry, there is basically a unique way to label the edges of the dual tiling in this way in a compatible fashion.) The boundary of every dual $4g$-gon now reads cyclically, counterclockwise, the relation $[a_{1},b_{1}]\cdots[a_{g},b_{g}]$. See Figure 5.1.
We pick one of the vertices to be the identity element and mark it by $o$. Any other vertex $v$ of $\text{Cay}(\Gamma)$ now corresponds to the element of $\Gamma$ represented by any of the paths from $o$ to $v$. The left action of $\Gamma$ on $\text{Cay}(\Gamma)$ extends to an action by isometries on $\mathbb{H}^{2}$. The isometry corresponding to each non-trivial element of $\Gamma$ is hyperbolic, which means that it admits a (unique, bi-infinite) geodesic axis in $\mathbb{H}^{2}$: the isometry acts on this axis by translation. These facts can be found, e.g., in [FM11, §1].
5.2 $\prod$-shaped paths for elements in $\Gamma$
The following definition will play a key role in the sequel. Roughly speaking, it is the appropriate replacement for surface groups of the fact that any element $\gamma$ of a free group can be expressed as $bhb^{-1}$ where $h$ is the cyclic reduction of $\gamma$.
Definition 5.1 ($\prod$-path).
Let $1\neq\gamma\in\Gamma$. Consider the unique (bi-infinite) geodesic $\rho\subset\mathbb{H}^{2}$ which is the axis of the isometry given by $\gamma$. The $\prod$-path corresponding to $\gamma$, which we denote by $\prod_{\gamma}$, is the path in $\mathbb{H}^{2}$ made out of three geodesics arcs: a geodesic arc $[o,x]$ from $o$ to $x\in\rho$ which is perpendicular to $\rho$, the geodesic arc $[x,\gamma.x]\subset\rho$, and the geodesic arc $[\gamma.x,\gamma.o]$ (which is, too, perpendicular to $\rho$).
See Figure 5.3 below. Note that for any hyperbolic isometry of $\mathbb{H}^{2}$, each half-space at either side of the axis is invariant under the isometry. This means that $o$ and $\gamma.o$ lie on the same side of the axis, so that $\prod_{\gamma}$ is made of three geodesic arcs with two right-turns, or two left-turns, between them (unless $o$ happens to lie on the axis of $\gamma$, in which case $\prod_{\gamma}$ is simply a geodesic arc).
We will need the following easy fact from hyperbolic geometry:
Fact 5.2.
Let $L$ be a geodesic in $\mathbb{H}^{2}$ and $L_{1},L_{2}$ be two geodesic rays perpendicular to $L$, emanating from two distinct points $y_{1}$ and $y_{2}$, respectively, on $L$, and at the same side of $L$. Denote $d_{1}=d(y_{1},y_{2})>0$. Let $z$ be a point on $L_{2}$ at distance $d_{2}\geq 0$ from $y_{2}$. Then the distance $d(z,L_{1})$ of $z$ from $L_{1}$ grows monotonically with each of $d_{1}$ and $d_{2}$, and for every fixed $d_{1}>0$,
$d(z,L_{1})\stackrel{{\scriptstyle d_{2}\to\infty}}{{\to}}\infty.$
[…]
$\scriptstyle d(z,L_{1})$
Figure 5.2: Illustration of Fact 5.2 (half-plane model of $\mathbb{H}^{2}$)
Fact 5.2 is illustrated in Figure 5.2. It follows immediately from the explicit formula $\sinh d(z,x)=\sinh(d_{1})\cosh(d_{2})$, that may be found in [Mar75, Thm. 32.21(a’)]. For completeness, we give an elementary argument.
Proof of Fact 5.2.
Assume without loss of generality that $L_{1}$ is the geodesic $x=0$ in the upper half plane model of $\mathbb{H}^{2}$. Then the distance between $z$ and $L_{1}$ is determined by the argument of $z$, and monotonically increases and tends to $\infty$ as the argument in $\left[0,\frac{\pi}{2}\right]$ decreases. This proves the claims when $d_{1}>0$ is fixed.
Now fix $d_{2}\geq 0$. Then $d(z,L_{1})$ is continuous as a function of $d_{1}$. Clearly $d(z,L_{1})\stackrel{{\scriptstyle d_{1}\searrow 0}}{{\to}}0$ and $d(z,L_{1})\stackrel{{\scriptstyle d_{1}\to\infty}}{{\to}}\infty$. So if it is not monotonically increasing, there must be two different values $\alpha_{1}<\alpha_{2}$ of $d_{1}$ with two corresponding points $z_{1}\neq z_{2}$ with $d(z_{1},L)=d(z_{2},L)$. Then $z_{1}$ must lie on the Euclidean interval from $z_{2}$ to 00. So the hyperbolic isometry $\Phi\colon t\mapsto ct$ maps $z_{1}$ to $z_{2}$ for some $c>1$. But this isometry shows that $z_{2}$ is at distance $d_{2}$ from the geodesic $\Phi(L)$ which lies strictly above $L$ — a contradiction. ∎
The following lemma is illustrated in Figure 5.3.
Lemma 5.3.
There is a constant $c_{1}=c_{1}(g)>0$ so that $\prod_{\gamma}$ and the geodesic arc $[o,\gamma.o]$ are each contained in a $c_{1}$-neighborhood of the other. Moreover, there are two points $z_{1},z_{2}\in[o,\gamma.o]$, with $z_{2}$ not closer to $o$ than $z_{1}$, such that the two corners $x$ and $\gamma.x$ of $\prod_{\gamma}$ satisfy $d_{\mathbb{H}^{2}}(x,z_{1})<c_{1}$ and $d_{\mathbb{H}^{2}}(\gamma.x,z_{2})<c_{1}$.
$o$$\gamma.o$$x$$\gamma.x$$\scriptstyle\mathrm{axis}(\gamma)$$\rho_{1}$$\rho_{3}$$\rho_{2}$$\prod_{\gamma}$$z_{1}$$z_{2}$
Figure 5.3: Illustration of Definition 5.1 and Lemma 5.3
Proof.
$\mathbb{H}^{2}$ has $\delta$-thin triangles in the sense of Rips and Gromov: every side of a geodesic triangle is contained in a $\delta$-neighborhood of the union of the other two sides [BH99, §III.H.1]. It follows that every side of a geodesic quadrilateral is contained in a $2\delta$-neighborhood of the union of the other three sides. Hence, $[o,\gamma.o]\subset{\cal N}_{2\delta}\left(\prod_{\gamma}\right)$.
Conversely, denote by $\rho_{1}=[o,x]$, $\rho_{2}=[x,\gamma.x]$ and $\rho_{3}=[\gamma.x,\gamma.o]$ the three geodesic arcs composing $\prod_{\gamma}$, so $x=\rho_{1}\cap\rho_{2}$ and $\gamma.x=\rho_{2}\cap\rho_{3}$. Note that $|\rho_{2}|$ is bounded from below by the injectivity radius of the genus-$g$ hyperbolic surface $\Gamma\backslash\mathbb{H}^{2}$. Hence, by Fact 5.2, there is some $c^{\prime}\geq 2\delta$ such that if $y\in\rho_{1}$ satisfies $d(y,x)>c^{\prime}$, then $d(y,\rho_{3})>4\delta$. If $\left|\rho_{1}\right|<2c^{\prime}$ we may simply take $z_{1}=o$ and $z_{2}=\gamma.o$ (recall that $\left|\rho_{3}\right|=\left|\rho_{1}\right|$). Otherwise, we reason as follows. By construction, the sets ${\cal N}_{2\delta}\left(\left\{y\in\rho_{1}:d(y,x)>c^{\prime}\right\}\right)$ and ${\cal N}_{2\delta}(\rho_{3})$ are disjoint. Thus $[o,\gamma.o]$ must visit a point $z_{1}\in{\cal N}_{2\delta}\left(\left\{y\in\rho_{1}:d(y,x)\leq 2c^{\prime}\right\}\right)$ before it reaches ${\cal N}_{2\delta}\left(\rho_{3}\right)$. Clearly, $d(z_{1},x)\leq 2c^{\prime}+2\delta$. Symmetrically, the inverse geodesic $[\gamma.o,o]$ must visit a point $z_{2}\in{\cal N}_{2\delta}\left(\left\{y\in\rho_{3}:d(y,\gamma.x)\leq 2c^{\prime}\right\}\right)$ before it reaches ${\cal N}_{2\delta}\left(\rho_{1}\right)$ and $d(z_{2},\gamma.x)\leq 2c^{\prime}+2\delta$. Finally,
$\rho_{1}\subset{\cal N}_{\delta}\left([o,z_{1}]\cup[z_{1},x]\right)\subseteq{\cal N}_{3\delta+2c^{\prime}}\left([o,z_{1}]\right),$
likewise $\rho_{3}\subset{\cal N}_{3\delta+2c^{\prime}}\left([z_{2},\gamma.o]\right)$, and
$\rho_{2}\subset{\cal N}_{2\delta}\left([x,z_{1}]\cup[z_{1},z_{2}]\cup[z_{2},\gamma.x]\right)\subseteq{\cal N}_{4\delta+2c^{\prime}}\left([z_{1},z_{2}]\right).$
The lemma is now proven with $c_{1}=4\delta+2c^{\prime}$. ∎
We now follow the terminology from [BS87]. Recall the tiling ${\cal T}$ of $\mathbb{H}^{2}$ that was introduced in §5.1. Denote by $hO$ the $4g$-gon with center $h.o$.
Definition 5.4 (Geodesic edge path [BS87, p. 455]).
Let $\rho$ be a finite-length, oriented geodesic arc in $\mathbb{H}^{2}$. Assume it begins in the interior of the $4g$-gon $h_{1}O$, and then visits $h_{2}O,h_{3}O$ and so on until it terminates at the interior of $h_{k}O$, and that it does not visit any vertex of the tiling ${\cal T}$ and does not coincide with any edge of ${\cal T}$. Then the corresponding geodesic edge path is the path $h_{1}.o-h_{2}.o-\ldots-h_{k}.o$ in $\text{Cay}(\Gamma)$. The corresponding geodesic word is the (reduced) word in $F(a_{1},\ldots,b_{g})$
$\left(h_{1}^{-1}h_{2}\right)\left(h_{2}^{-1}h_{3}\right)\ldots\left(h_{k-1}^{-1}h_{k}\right)$
(recall that if $hO$ and $h^{\prime}O$ share a common side, then the edge in $\text{Cay}(\Gamma)$ from $hO$ to $h^{\prime}O$ is labeled by $h^{-1}h^{\prime}\in\left\{a_{1}^{\pm 1},\ldots,b_{g}^{\pm 1}\right\}$). If $\rho$ goes through a vertex of ${\cal T}$ or coincides with an edge of ${\cal T}$, we deform it slightly as in [Ibid., Fig. 1], and then define the geodesic edge path as above using the deformed arc. [^11] To avoid a vertex, the deformation consists of taking a detour on either side of the vertex. To avoid an edge, it consists of moving the entire geodesic arc to either side of the edge. The geodesic arcs we use here never coincide with edges of ${\cal T}$. If $\rho$ needs to be deformed, the resulting geodesic edge path is not unique, but this is immaterial for our purposes.
Using geodesic edge paths, we have an analog of Lemma 5.3 in $\text{Cay}(\Gamma)$. It basically says that in $\text{Cay}(\Gamma)$ there are, too, a geodesic path from $o$ to $\gamma.o$ and a $\prod_{\gamma}$-like path which are close to each other. We focus on proper powers and on the precise properties that will be needed below. We view $\text{Cay}(\Gamma)$ as a metric space with edge-length $1$, and a geodesic path in it is any shortest path between two vertices (here and in the following lemma, vertices are vertices of $\text{Cay}(\Gamma)$ and not of the tiling, unless stated otherwise).
Lemma 5.5.
There is a constant $c_{2}=c_{2}(g)>0$ so that the following holds. Let $k\in\mathbf{Z}_{\geq 2}$ and $1\neq\gamma\in\Gamma$ be a proper $k^{\text{th}}$-power. Then there are $b,h\in\Gamma$ with $\gamma=bh^{k}b^{-1}$, a geodesic path $\tau_{\gamma}\subset\text{Cay}(\Gamma)$ from $o$ to $\gamma.o$, and four vertices $v_{0},v_{1},v_{2},v_{k}\in\tau_{\gamma}$, with $v_{j}$ not closer to $o$ than $v_{i}$ when $i<j$, such that $d_{\text{Cay}(\Gamma)}(bh^{j}.o,v_{j})<c_{2}$ for all $j=0,1,2,k$.
Proof.
Let $\gamma_{0}\in\Gamma$ with $\gamma=\gamma_{0}^{\,k}$. Consider the geodesic arc $\left[o,\gamma.o\right]$ in $\mathbb{H}^{2}$, (deform it a bit if need be), and let $\tau_{\gamma}$ be the associated geodesic edge path as in Definition 5.4. By [BS87, Thm. 2.8(b)], every geodesic edge path, and $\tau_{\gamma}$ in particular, is a geodesic in $\text{Cay}(\Gamma)$.
[…]
Denote by $D>0$ the diameter of the $4g$-gons. Every point on a geodesic arc in $\mathbb{H}^{2}$ lies in the $D$-neighborhood of its corresponding geodesic edge path, and even in the $D$-neighborhood of the vertices of the geodesic edge path. So we may find vertices $v_{0},v_{1},v_{2},v_{k}\in\tau_{\gamma}$ with $d_{\mathbb{H}^{2}}(z_{j},v_{j})<D$ for $j=0,1,2,k$. It is clear, by the definition of a geodesic edge path, that we may choose the $v_{j}$’s so that $v_{j}$ is not closer than $v_{i}$ to $o$ when $i<j$. As $d_{\mathbb{H}^{2}}(\gamma_{0}^{~{}j}.x,bh^{j}.o)<D$ for all $j=0,1,2,k$, we obtain that
$d_{\mathbb{H}^{2}}\left(bh^{j}.o,v_{j}\right)<4\delta+c_{1}\left(g\right)+2D.$
Finally, by the S̆varc-Milnor Lemma (e.g., [BH99, Prop. I.8.19]), the embedding above of $\text{Cay}(\Gamma)$ in $\mathbb{H}^{2}$ is a quasi-isometry, and bounded distances remain bounded. In particular, there is some $c_{2}\left(g\right)>0$ so that $d_{\text{Cay}(\Gamma)}\left(bh^{j}.o,v_{j}\right)<c_{2}\left(g\right)$ for $j=0,1,2,k$. ∎
The following lemma considers words with letters in an arbitrary (finite) subset ${\cal S}\subset\Gamma$. Let $w$ be such a word of length $p$. We let $w_{[t_{1},t_{2})}$ denote the subword consisting of the letters at positions $t_{1},t_{1}+1,\ldots,t_{2}-1$ (here $11\leq t_{1}\leq t_{2}\leq p+$ 1), and $w_{[t_{1},t_{2}]}$ the subword consisting of the letters at positions $t_{1},\ldots,t_{2}$ (here $1\leq t_{1}\leq t_{2}\leq p$).
Lemma 5.6.
There is a constant $c_{3}=c_{3}(g)>0$ so that the following holds. Fix any finite subset $\mathcal{S}\subset\Gamma$ so that $\max_{\gamma\in\mathcal{S}}|\gamma|\leq d$, and let $w$ be any not-necesssarily-reduced word of length $p$ with letters in $\mathcal{S}$ representing a non-trivial $k^{\text{th}}$ power in $\Gamma$ for some $k\geq 2$. Then there are $11\leq t_{1}\leq t_{2}\leq t_{3}\leq t_{4}\leq p+$ 1, and elements $b,h,u_{1},\ldots,u_{4}\in\Gamma$ with $\left|u_{i}\right|\leq c_{3}\left(d+\log p\right)$, such that
$\left(w_{[1,t_{1})}u_{1}\right)\cdot\left(u_{1}^{-1}w_{[t_{1},t_{2})}u_{2}\right)\cdot\left(u_{2}^{-1}w_{[t_{2},t_{3})}u_{3}\right)\cdot\left(u_{3}^{-1}w_{[t_{3},t_{4})}u_{4}\right)\cdot\left(u_{4}^{-1}w_{\left[t_{4},p\right]}\right)$
is of the form $b\cdot h\cdot h\cdot h^{k-2}\cdot b^{-1}$. Namely,
$w_{[1,t_{1})}u_{1}\equiv_{\Gamma}b,\qquad$ $u_{1}^{-1}w_{[t_{1},t_{2})}u_{2}\equiv_{\Gamma}u_{2}^{-1}w_{[t_{2},t_{3})}u_{3}\equiv_{\Gamma}h$
$u_{3}^{-1}w_{[t_{3},t_{4})}u_{4}\equiv_{\Gamma}h^{k-2},\qquad$ $u_{4}^{-1}w_{\left[t_{4},p\right]}\equiv_{\Gamma}b^{-1}.$
Here $\equiv_{\Gamma}$ denotes equality in $\Gamma$.
Proof.
All the distances in this proof are measured in $\text{Cay}(\Gamma)$. Let $1\neq\gamma\in\Gamma$ denote the element represented by $w$, and let $b,h\in\Gamma$, $\tau_{\gamma}$ and $v_{0},v_{1},v_{2},v_{k}\in\tau_{\gamma}$ be the elements of $\Gamma$, the geodesic and the vertices given by Lemma 5.5. We may expand $w$ to a word of length at most $dp$ in the generators $\left\{a_{1}^{\pm},\ldots,b_{g}^{\pm}\right\}$ of $\Gamma$, and the expanded word then represents a path $\mathfrak{p}_{w}$ of length at most $dp$ from $o$ to $\gamma.o$. The Cayley graph $\text{Cay}(\Gamma)$ is a hyperbolic space (in the sense of Rips-Gromov) with $\delta$-thin triangles for some $\delta>0$. A standard fact in hyperbolic geometry (e.g., [BH99, Prop. III.H.1.6]) gives that as $\tau_{\gamma}$ is a geodesic in $\text{Cay}(\Gamma)$ between $o$ and $\gamma.o$, every point $x\in\tau_{\gamma}$ is at distance at most $\delta\left|\log_{2}\left|\mathfrak{p}_{w}\right|\right|+1\leq\delta\left|\log_{2}(pd)\right|+1$ from $\mathfrak{p}_{w}$. Together with Lemma 5.5, this shows that each of the points $b.o,bh.o,bh^{2}.o$ and $bh^{k}.o$ are rather close to $\mathfrak{p}_{w}$. It remains to show we can find four points along $\mathfrak{p}_{w}$ which are close to $b.o,bh.o,bh^{2}.o,bh^{k}.o$ and ordered correctly.
First, as every vertex on $\mathfrak{p}_{w}$ is at distance at most $d$ from the beginning of a letter of $w$, there is a vertex $y_{0}$ on $\mathfrak{p}_{w}$ which is at the beginning of a letter of $w$ and with $d(v_{0},y_{0})<\delta\left|\log_{2}\left(pd\right)\right|+1+d$, so
$d(b.o,y_{0})\leq d(b.o,v_{0})+d(v_{0},y_{0})<c_{2}+\delta\left|\log_{2}\left(pd\right)\right|+1+d<c^{\prime}\left(d+\log p\right),$
for some $c^{\prime}>0$ depending only on $g$. Set $u_{1}$ to be a word representing a shortest path in $\text{Cay}(\Gamma)$ from $y_{0}$ to $b.o$, and let $t_{1}$ be the smallest integer so that $w_{[1,t_{1})}$ ends at $y_{0}$. Now consider the path from $v_{0}$ to $\gamma.o$ given by $\mathfrak{q}\cdot w_{[t_{1},p]}$, where $\mathfrak{q}$ is a geodesic path from $v_{0}$ to $y_{0}$ (so $\left|\mathfrak{q}\right|<\delta\left|\log_{2}(pd)\right|+1+d$). As $v_{1}$ lies on a geodesic from $v_{0}$ to $\gamma.o$, there is a vertex $y_{1}$ on $\mathfrak{p}_{w}$, at the beginning of a letter of $w_{[t_{1},p]}$, such that
$d(v_{1},y_{1})<\delta\left|\log_{2}\left(d(p+1-t_{1})+\left|\mathfrak{q}\right|\right)\right|+1+d+\left|\mathfrak{q}\right|,$
so $d(bh.o,y_{1})<c^{\prime\prime}(d+\log p)$ for some $c^{\prime\prime}>0$ depending only on $g$. We let $u_{2}$ be a word representing a shortest path from $y_{1}$ to $bh.o$ and $t_{2}$ the smallest integer so that $t_{2}\geq t_{1}$ and $w_{[1,t_{2})}$ ends at $y_{1}$.
In the same manner, and enlarging the constant $c^{\prime\prime}$ as needed at each step, we find points $y_{2}$ and $y_{k}$ on the path of $w$, words $u_{3}$ and $u_{4}$ and times $t_{3}$ and $t_{4}$ so that the statement of the lemma holds. ∎
5.3 Proof of Theorem 1.8
As is explained in [MdlS24b, §6], the proof of Theorem 1.8 can be reduced to bounding the exponential growth rate of the probability that a product $\boldsymbol{\gamma}_{1}\cdots\boldsymbol{\gamma}_{p}$ of i.i.d. finitely supported random group elements $\boldsymbol{\gamma}_{i}\in\Gamma$ (i.e., a random walk on $\Gamma$ with a general finitely supported generating measure) lands on a proper power.
In the case of the free group, the requisite bound follows by a simple spectral argument that dates back to [Fri03, Lem. 2.4]. This argument can be adapted to the setting of surface groups due to Lemma 5.6.
Proof of Theorem 1.8.
Fix a self-adjoint $x=\sum_{\gamma\in\Gamma}\alpha_{\gamma}\gamma$ with $|x|=d$. By [MdlS24b, Prop. 6.3] and as surface groups satisfy the rapid decay property [Jol90, Thm. 3.2.1], we can assume that $x$ has positive coefficients, and by homogeneity we can assume the coefficients sum to one. Thus we can write $x^{p}=\mathbb{E}[\boldsymbol{\gamma}_{1}\cdots\boldsymbol{\gamma}_{p}]$ for every $p\in\mathbf{N}$, where $\boldsymbol{\gamma}_{i}$ are i.i.d. random elements of $\Gamma$ with $\mathbb{P}\left[\boldsymbol{\gamma}_{i}=\gamma\right]=\alpha_{\gamma}$. In particular, (1.3) yields
$u_{1}\left(x^{p}\right)=-\tau{\left(\lambda(x)^{p}\right)}+\sum_{k=2}^{pd}\left(\omega(k)-1\right)\sum_{v\in\Gamma_{{\rm np}}}\mathbb{E}\left[1_{\boldsymbol{\gamma}_{1}\cdots\boldsymbol{\gamma}_{p}\equiv_{\Gamma}v^{k}}\right],$
where $\Gamma_{\rm np}$ denotes the non-powers in $\Gamma$. Here we used that $\boldsymbol{\gamma}_{1}\cdots\boldsymbol{\gamma}_{p}\equiv_{\Gamma}v^{k}$ implies that $k\leq|v^{k}|\leq|\boldsymbol{\gamma}_{1}|+\cdots+|\boldsymbol{\gamma}_{p}|\leq pd$, and that $\mathbb{E}[1_{\boldsymbol{\gamma}_{1}\cdots\boldsymbol{\gamma}_{p}\equiv_{\Gamma}1}]=\mathbb{E}[\tau(\lambda(\boldsymbol{\gamma}_{1}\cdots\boldsymbol{\gamma}_{p}))]=\tau(\lambda(x)^{p})$.
By Lemma 5.6, we have for every $k\geq 2$ and $\gamma_{1},\ldots,\gamma_{p}\in\Gamma$
$\sum_{v\in\Gamma_{\text{np}}}\mbox{$1_{\gamma_{1}\cdots\gamma_{p}\equiv_{\Gamma}v^{k}}$}$$\leq$ $\sum_{1\leq t_{1}\leq\cdots\leq t_{4}\leq p+1}\sum_{\begin{subarray}{c}b,h,u_{1},\ldots,u_{4}\in\Gamma\\ |u_{i}|\leq c_{3}(d+\log p)\end{subarray}}1_{\gamma_{1}\cdots\gamma_{t_{1}-1}\equiv_{\Gamma}bu_{1}^{-1}}1_{\gamma_{t_{1}}\cdots\gamma_{t_{2}-1}\equiv_{\Gamma}u_{1}hu_{2}^{-1}}\times$
[…]
As
$1_{\gamma_{1}\cdots\gamma_{p}\equiv_{\Gamma}\gamma}=\left\langle\delta_{\gamma},\lambda(\gamma_{1}\cdots\gamma_{p})\delta_{e}\right\rangle,$
this yields
$\sum_{v\in\Gamma_{\text{np}}}\mbox{$\mathbb{E}\left[1_{\boldsymbol{\gamma}_{1}\cdots\boldsymbol{\gamma}_{p}\equiv_{\Gamma}v^{k}}\right]$}\leq\sum_{1\leq t_{1}\leq\cdots\leq t_{4}\leq p+1}\sum_{\begin{subarray}{c}b,h,u_{1},\ldots,u_{4}\in\Gamma\\ |u_{i}|\leq c_{3}(d+\log p)\end{subarray}}\left\langle\delta_{bu_{1}^{-1}},\lambda(x)^{t_{1}-1}\delta_{e}\right\rangle\times\\ \left\langle\delta_{u_{1}hu_{2}^{-1}},\lambda(x)^{t_{2}-t_{1}}\delta_{e}\right\rangle~{}\left\langle\delta_{u_{2}hu_{3}^{-1}},\lambda(x)^{t_{3}-t_{2}}\delta_{e}\right\rangle~{}\left\|\lambda(x)\right\|^{t_{4}-t_{3}}~{}\left\langle\delta_{u_{4}b^{-1}},\lambda(x)^{p+1-t_{4}}\delta_{e}\right\rangle.$
Now note that the sum of $t_{1},\ldots,t_{4}$ has at most $(p+1)^{4}$ terms, while the sum over $u_{1},\ldots,u_{4}$ has at most $(4g+1)^{4c_{3}(d+\log p)}$ terms. Moreover, as
$\sum_{v\in\Gamma}|\langle\delta_{v},\lambda(x)^{t}\delta_{e}\rangle|^{2}=\|\lambda(x)^{t}\delta_{e}\|^{2}\leq\|\lambda(x)\|^{2t},$
we can apply Cauchy-Schwarz to the remaining sums over $b,h$ to estimate
$u_{1}(x^{p})$ $\leq|\tau{\left(\lambda(x)^{p}\right)}|+(pd)^{2}\max_{2\leq k\leq pd}\sum_{v\in\Gamma_{{\rm np}}}\mathbb{E}\big{[}1_{\boldsymbol{\gamma}_{1}\cdots\boldsymbol{\gamma}_{p}\equiv v^{k}}\big{]}$
$\leq\left[1+(pd)^{2}(p+1)^{4}(4g+1)^{4c_{3}(d+\log p)}\right]\|\lambda(x)\|^{p}.$
The conclusion follows directly. ∎
6 Proof of Theorem 1.1
Before we complete the proof, we must dispense with a cosmetic issue. Let
$\pi_{n}=\mathrm{std}\circ\phi_{n},$
where $\phi_{n}$ is a uniformly distributed random element of $\mathrm{Hom}(\Gamma_{g},S_{n})$. In Theorem 1.7, we have proved an effective asymptptic expansion of $\mathbb{E}[\mathop{\mathrm{Tr}}\pi_{n}(\gamma)]$ in powers of $\frac{1}{n}$. However, since $\pi_{n}:\Gamma\to\mathcal{U}(n-1)$ is a representation of dimension $n-1$, this differs slightly from the setting of section 1.1 where $\pi_{n}$ is $n$-dimensional.
This minor discrepancy is readily resolved by noting that the statement and proof of Theorem 1.5 extend verbatim to the case that $\pi_{n}$ is $(n-1)$-dimensional, provided that we still understand the normalized trace to be defined by
$\mathop{\mathrm{tr}}\pi_{n}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{1}{n}\mathop{\mathrm{Tr}}\pi_{n}.$
We will therefore impose this convention in the remainder of this section. [^12] The issue is purely cosmetic in nature, since the proof of Theorem 1.7 could also be readily adapted to yield an effective asymptotic expansion in powers of $\frac{1}{n-1}$.
We can now proceed to assemble all the pieces of the proof.
Proof of Theorem 1.1.
The effective asymptotic expansion of Assumption 1.3 follows immediately from Theorem 1.7 and (1.1) with
$u_{1}(\gamma)=a_{0}(\gamma)-1,\qquad\quad u_{k}(\gamma)=a_{k-1}(\gamma)\quad\text{for}\quad k\neq 1.$
Moreover, it is shown in [MP23, Thm. 1.2] that $a_{-1}(\gamma)=1_{\gamma=1}$ and $a_{0}(\gamma)=\omega(\gamma)$. Thus
$u_{0}(\gamma)=1_{\gamma=1}=\tau(\lambda(\gamma)),\qquad\qquad u_{1}(\gamma)=\omega(\gamma)-1.$
Thus Assumption 1.3 holds, and Assumption 1.4 follows from Theorem 1.8. Consequently, Theorem 1.5 yields the strong convergence upper bound
$\|\pi_{n}(x)\|\leq\|\lambda(x)\|+o(1)\quad\text{with probability}\quad 1-o(1)$
as $n\to\infty$ for every $x\in\mathbf{C}[\Gamma]$. As surface groups have the unique trace property [dlH88], the corresponding strong convergence lower bound
$\|\pi_{n}(x)\|\geq\|\lambda(x)\|-o(1)\quad\text{with probability}\quad 1-o(1)$
follows from the upper bound, see, e.g., [MdlS24b, §5.3]. ∎